### v1

# ACING: Actor-Critic for Instruction Learning in Black-Box Large Language Models

Abstract
The effectiveness of Large Language Models (LLMs) in solving tasks vastly depends on the quality of the instructions, which often require fine-tuning through extensive human effort. This highlights the need for automated instruction optimization; however, this optimization is particularly challenging when dealing with black-box LLMs, where model parameters and gradients remain inaccessible.
We propose ACING, a task-specific prompt optimization approach framed as a stateless continuous-action Reinforcement Learning (RL) problem, known as the continuum bandit setting. ACING leverages an actor-critic-based method to optimize prompts, learning from non-differentiable reward signals. We validate ACING by optimizing prompts for ChatGPT on 30 instruction-based tasks, as well as a summarization task. ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points compared to the best baseline considered. Furthermore, ACING not only recovers but also surpasses human-crafted expert instructions, achieving up to a 39 percentage point improvement over human benchmarks.
ACING: Actor-Critic for Instruction Learning
in Black-Box Large Language Models
[…]
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of tasks Zhao et al. (2024); Touvron et al. (2023). This success is largely attributed to their strong instruction-following capabilities, which enable adaptation to diverse downstream applications Chen et al. (2023); Liu et al. (2023). These instructions, commonly referred to as prompts, play a crucial role in the performance of LLMs Wei et al. (2022); Zhu et al. (2024); Liu et al. (2023). Given the importance of prompts, researchers are increasingly interested in automating their optimization to reduce the need for manual adjustments, which is often a labor-intensive and costly process Reynolds and McDonell (2021); Mishra et al. (2021). As a result, developing efficient methods for automatic prompt optimization has become highly important to maximize LLMs performance.
Several methods for automated prompt optimization have been proposed in the literature. Some focus on optimizing continuous prompts, known as soft prompts Zhong et al. (2021); Li and Liang (2021), which are typically fed into a LLM after its embedding layer since they do not represent meaningful language that the LLM can interpret directly. Other methods aim at optimizing discrete prompts, known as hard prompts, by either generating diverse prompts Zhou et al. (2023) or refining existing ones Pryzant et al. (2023). For both soft and hard prompts, some approaches rely on gradient-based methods Shin et al. (2020); Lester et al. (2021); Li and Liang (2021). However, these methods require access to the gradients of the LLMs and are therefore limited to white-box LLMs. Meanwhile, the most powerful LLMs today are typically black-box models (e.g., ChatGPT OpenAI (2023a) and GPT-4 OpenAI (2023b)).
To overcome these challenges, recent studies have introduced gradient-free techniques for optimizing prompts in black-box LLMs Zhou et al. (2023); Prasad et al. (2023); Pryzant et al. (2023). While promising, these approaches rely on heuristic local search methods and fail to fully utilize the observation history, including previously tested instructions and their outcomes, when selecting new instructions to test. Consequently, they are less effective at balancing the inherent exploration-exploitation trade-off, leading to inefficient querying. This inefficiency becomes problematic and impractical when API calls to black-box LLMs involve significant time and financial costs.
A recent line of work has focused on optimizing instructions for black-box LLMs by searching through soft prompts and transforming them into discrete prompts using white-box LLMs. InstructZero Chen et al. (2024) optimizes discrete prompts for ChatGPT by applying Bayesian optimization over soft prompts. Building on the same line, INSTINCT Lin et al. (2024) replaces Bayesian optimization with NeuralUCB Zhou et al. (2020), a critic algorithm, to select actions, from a finite pool of soft prompts, initially generated with Sobol Eriksson et al. (2019).
Inspired by the success of actor-critic methods for control problems Haarnoja et al. (2018, 2019) and generative adversarial networks Goodfellow et al. (2014), our approach combines the hybrid white-box/black-box framework with an actor-critic Reinforcement Learning (RL) algorithm, operating in a stateless, infinite continuous-action space, and constrained by a strict budget of API calls. The actor aims to maximize the critic’s value while remaining as stochastic as possible to enhance exploration and simultaneously exploit promising actions. This approach increases the likelihood of finding optimal prompts and enables more efficient use of the limited number of API calls. ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points. Furthermore, our optimized prompts not only recover but also surpass human-crafted expert instructions, achieving up to a 39 percentage point improvement against human benchmarks.
An implementation of ACING is available at https://github.com/salmakh1/ACING.
2 Problem Formulation
2.1 Prompt Learning Objective
We aim to improve the performance of a black-box LLM, denoted by $f$, which can only be accessed through its API, while its internal parameters remain unknown. Given a task represented by an (unknown) distribution $(x,y)\sim\mathcal{D}$—where $x$ denotes possible inputs and $y$ the corresponding correct outputs—our goal is to find the optimal prompt $\tau^{\star}$ that maximizes the likelihood of $f$ producing correct outputs for a given task. This is evaluated using a scoring function $q(\cdot,\cdot)\in[0,1]$. The black-box model $f$ processes an input formed by concatenating ($\oplus$) the prompt $\tau$ with the sentence $x$, producing a predicted output $\hat{y}=f(\tau\oplus x)$. More formally, the objective is to maximize the expected score of the LLM in solving the task $\mathcal{D}$, defined as:
$\max_{\tau}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau\oplus x))\right].$ (1)
To achieve this, we utilize a validation dataset $\mathcal{V}=\{(x_{j},y_{j})\}_{j=1}^{m}$, where each pair consists of an input sentence $x_{j}$ and its corresponding ground truth output $y_{j}$. Our objective is to find the prompt that maximizes the scoring function $q(\cdot,\cdot)$ across the validation dataset, where $q(\hat{y}_{j},y_{j})$ measures the quality of the predicted output $\hat{y}_{j}$ against the true output $y_{j}$. Thus, our objective becomes finding the prompt $\tau^{\star}$ that maximizes the average score over the validation set $\mathcal{V}$. The derived prompt $\tau^{\star}$ is then evaluated on a separate test set $\mathcal{T}=\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{m^{\prime}}$ to assess its generalization performance.
2.2 Learning over a Continuous Space
Directly optimizing the prompt $\tau$ for the black-box LLM model $f$ presents substantial challenges due to the discrete combinatorial nature of token selection in $\tau$. To mitigate this challenge, prior approaches like InstructZero Chen et al. (2024) and INSTINCT Lin et al. (2024) employ a white-box LLM, represented by $h$, to convert the discrete optimization problem into a continuous one by introducing a soft prompt vector $\mathbf{z}\in\mathbb{R}^{d}$, which is a continuous $d$-dimensional vector representing the token embedding of a set of virtual tokens.
The white-box model $h$ has access to a dataset of exemplars, =$\mathcal{E}=\{(u_{j},v_{j})\}_{j=1}^{k}$, where each pair $(u_{j},v_{j})$ defines input-output text sequences that exemplify a downstream task. The white-box LLM $h$ serves as a proxy, mapping the soft prompt $\mathbf{z}$ into a discrete prompt $\tau$ suitable for input to the black-box LLM. Specifically, given the exemplars $\mathcal{E}$ and the vector $\mathbf{z}$, their concatenation is input to the white-box LLM, generating the discrete prompt $\tau(\mathbf{z})=h(\mathbf{z},\mathcal{E})$. This generated prompt $\tau(\mathbf{z})$ is prepended to a test input $x_{j}$ from the validation set $\mathcal{V}$, and the combined input is provided to the black-box LLM $f$ to generate an output $\hat{y}_{j}=f(\tau(\mathbf{z})\oplus x_{j})$. The output $\hat{y}_{j}$ is then evaluated against the true output $y_{j}$ using the scoring function $q(\hat{y}_{j},y_{j})$. By using a fixed set of exemplars $\mathcal{E}$, the original discrete problem (Equation (1)) of finding the optimal prompt $\tau$ is effectively transformed into a continuous optimization problem over the soft prompt vector $\mathbf{z}$, as follows:
$\max_{\mathbf{z}\in\mathbb{R}^{d}}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau(\mathbf{z})\oplus x))\right].$ (2)
2.3 Reducing the Continuous Space
The soft prompt $\mathbf{z}$ is typically high-dimensional (e.g., $d=5120\times N_{z}$, with $N_{z}$ the number of virtual tokens, when $h$ is Vicuna-13B Chiang et al. (2023)), which makes it a challenging problem. Therefore, we employ random projection techniques to reduce the input dimension as done in prior works Chen et al. (2024); Lin et al. (2024). Specifically, given a matrix $P\in\mathbb{R}^{d\times d^{\prime}}$ where $d^{\prime}\ll d$, with randomly sampled elements, and a continuous vector $\mathbf{a}\in\mathbb{R}^{d^{\prime}}$, the vector $\mathbf{z}=P\mathbf{a}$ is used as the soft prompt. By substituting $\mathbf{z}$ with $P\mathbf{a}$, the input variable to be optimized in Eq. 2 becomes $\mathbf{a}$, thereby reducing the input dimension of the optimization problem to $d^{\prime}$:
$\max_{\mathbf{a}\in\mathbb{R}^{d^{{}^{\prime}}}}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau(P\mathbf{a})\oplus x))\right].$ (3)
For fairness with previous works, the reduced input dimension $d^{\prime}$, referred to as the intrinsic dimension, is chosen as $d^{\prime}=10$ Chen et al. (2024); Lin et al. (2024). An ablation study on the intrinsic dimension is discussed in § 5.4.3. Furthermore, to maintain consistency with these prior studies, we similarly use a projection matrix $P$ with elements uniformly and independently sampled from $\text{Uniform}(-1,1)$, and bound the vector such that $\mathbf{a}\in[0,1]^{d^{\prime}}$.
3 Actor-Critic Approach for Instruction Learning in Black-Box LLMs
3.1 Instruction Learning as an RL Problem
In this work, we frame the prompt-learning problem as an agent navigating an infinite action space within a $d^{{}^{\prime}}$-dimensional unit hypercube, searching within an infinite number of actions $\mathbf{a}\in[0,1]^{d^{{}^{\prime}}}$, in contrast to previous works that consider a finite number of actions initially generated with Sobol Eriksson et al. (2019); Lin et al. (2024). As part of the environment, these continuous actions are projected into a higher-dimensional space (§ 2.3) and then transformed into discrete prompts, through a white-box LLM (§ 4). These discrete prompts are evaluated by a black-box LLM over a validation dataset, after which a reward is computed and returned to the agent. We consider a finite horizon $T$, representing the maximum number of actions played (i.e., the budget of evaluated instructions), after which the agent ceases exploration. The agent’s goal is to discover the optimal continuous action, $\mathbf{a}^{\star}$, which produces the optimal discrete prompt and the highest reward.
Unlike in general RL settings, where the agent encounters multiple states with varying reward distributions, the agent we consider operates in a stateless, stochastic environment, commonly referred to as a stochastic bandit setting Slivkins (2019); Lattimore and Szepesvári (2020). In this context, at each iteration $t\leq T$ (where $t$ is incremented with every black-box LLM call), the agent’s action $\mathbf{a}_{t}$ generate stochastic rewards, $r_{t}\in[0,1]$, without affecting the reward distributions associated with those actions. The prompt-learning objective is modeled as a stochastic best-arm identification problem in a stateless, continuous action space RL, specifically in an infinite continuum bandit setting.
Algorithm 1 Stateless Soft Actor-Critic
0:  Evaluation budget $T$
1:  Initialize $\boldsymbol{\theta}$, $\mathbf{w}$ (e.g., randomly)
2:  for $t\leftarrow 1$ to $T$ do
3:     Choose action $\mathbf{a}_{t}\sim\pi(\cdot\mid\boldsymbol{\theta})$
4:     Take action $\mathbf{a}_{t}$ and observe reward $r_{t}$
5:     $\mathbf{w}\leftarrow\mathbf{w}-\lambda\hat{\nabla}\mathbf{J}_{Q}(\mathbf{w})\quad$ (Equation (5))
6:     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\beta\hat{\nabla}\mathbf{J}_{\pi}(\boldsymbol{\theta})\quad$ (Equation (6))
7:     $\alpha\leftarrow\alpha-\gamma\hat{\nabla}\mathbf{J}_{\alpha}(\alpha)\quad$ (Equation (7))
8:  end for
3.2 Actor-Critic Continuum Bandits
RL techniques are generally used to learn optimal policies and actions leading to best rewards. These techniques can be classified into value-based (critic-centric), policy-gradient (actor-centric), or hybrid approaches. Actor-critic algorithms merge the benefits of both by employing two key components: the actor, responsible for policy optimization, and the critic, which evaluates the policy (Sutton and Barto, 2018). This hybrid approach has demonstrated success in various domains, including continuous control problems Mnih et al. (2016); Haarnoja et al. (2018, 2019) and with generative adversarial networks (GANs) (Goodfellow et al., 2014), where a generator and a critic network collaboratively optimize performance.
Figure 1: Pipeline of ACING. In each iteration, a soft prompt along with several examples of the target task are provided to the white-box LLM to generate an instruction. This instruction is then used to query the black-box LLM, which produces answers to the target task queries. The resulting score is returned to the agent as a reward, which is used to update its networks and adjust its policy. Both LLMs remain frozen throughout the process.
Inspired by the success of actor-critic approaches, we introduce a (stateless) actor-critic algorithm, as provided in Algorithm 1, tailored to our infinite continuum bandit setting, enabling autonomous learning of effective prompts with a constrained evaluation budget and outperforming previous state-of-the-art black-box prompt learning approaches. Both the actor and the critic are parameterized by neural networks. In typical RL problems, the agent interacts with an environment with varying states, so its policy depends on the current state. That is, the policy network generates actions based on the current state $s$, $\pi(.|s;\theta)$, where $\theta$ denotes the parameters. However, in the problem we consider, the setting is stateless. Therefore, we define a policy network $\pi(.;\theta)$ with a constant input that produces a $d^{{}^{\prime}}$-dimensional vector as output $\mathbf{a}$, representing the continuous action. For the critic, we use a neural network with a $d^{{}^{\prime}}$-dimensional input (the action) and a single output that estimates the action’s quality, thereby assessing the quality of the policy that proposed it. The critic network’s parameters are trained to minimize the following:
$\min_{\mathbf{w}}J_{Q}(\mathbf{w})\triangleq\mathbb{E}_{\mathbf{a}\sim\mathcal{D}}\left[\frac{1}{2}\left(Q_{\mathbf{w}}\left(\mathbf{a}\right)-r\left(\mathbf{a}\right)\right)^{2}\right],$ (4)
which can be optimized with stochastic gradients
^$\hat{\nabla}_{\theta}J_{Q}(\mathbf{w})=\nabla_{\mathbf{w}}Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)\left(Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)-r\left(\mathbf{a}_{t}\right)\right),$ (5)
where $Q$ is the critic-network with parameters $\mathbf{w}$. The actor-critic learning process alternates between policy-network and critic-network updates.
3.3 Enhancing Exploration with Entropy
In general, RL learning approaches aim to maximize the expected cumulative rewards up to a horizon $T$, expressed as $\max_{\pi(\cdot)}\sum_{t=1}^{T}\mathbb{E}_{a_{t}\sim\pi(\cdot)}\left[r(a_{t})\right].$ In contrast, black-box prompt learning, framed as a best-arm identification problem, seeks the best prompt without necessarily maximizing the expected cumulative rewards. To maximize the cumulative rewards, the agent must manage the exploration-exploitation trade-off. Conversely, for the best-arm identification problem, the agent should explore as extensively as possible within the $T$ evaluation budget to potentially identify the best action, leading to a pure exploration setting. Although it might seem that the agent should act purely at random to find the best action, some degree of exploitation may still be beneficial to guide the agent toward the optimal choice.
To achieve a high exploration rate while exploiting prior data, we consider the maximum entropy objective, as discussed in previous RL works Ziebart (2010); Haarnoja et al. (2018). This objective promotes stochastic policies by augmenting the reward with an entropy term, expressed as:
$\min_{\theta}J_{\pi}(\theta)\triangleq\mathbb{E}_{\mathbf{a}\sim\pi(.;\theta)}\left[\alpha\log\left(\pi\left(\mathbf{a};\theta\right)\right)-Q_{\mathbf{w}}\left(\mathbf{a})\right)\right],$ (6)
where the temperature $\alpha$ controls the stochasticity of the agent. The above objective is a simplification of the objective in the Soft Actor-Critic (SAC) algorithm Haarnoja et al. (2018), eliminating the state, as we consider a stateless environment and focus on instantaneous rather than cumulative rewards. The term $-\mathbb{E}_{a\sim\pi(.;\theta)}\left[\log\pi(a)\right]$ represents the entropy of the policy, which we seek to maximize alongside the critic values, as entropy maximization is crucial for encouraging the actor to explore different actions. A high $\alpha$ encourages exploration, whereas a lower value of $\alpha$ favors exploitation.
Furthermore, to avoid fine-tuning, and to achieve adaptive exploration, the entropy coefficient $\alpha$ can be adjusted dynamically to maintain a target entropy level Haarnoja et al. (2019), $H_{\text{target}}$, ensuring sufficient exploration, as follows:
$\min_{\alpha}\mathbf{J}_{\alpha}\triangleq-\mathbb{E}_{a\sim\pi(.)}\left[\alpha\cdot\left(\log\pi(\mathbf{a};\theta)+H_{\text{target}}\right)\right].$ (7)
We use a stochastic approximation of gradients, specifically the Adam algorithm Kingma and Ba (2017); Reddi et al. (2018), to minimize the losses via stochastic gradient descent. The learning rates $\lambda$, $\beta$, and $\gamma$ are used for the Q-network, the policy networks, and the temperature, respectively.
Figure 2: Illustration of the prompt generation and testing inside the environment.
4 Methodology
The methodology of our proposed algorithm, ACING, is illustrated in Fig. 1, showing the actor-critic interaction with the environment. In Fig. 2, the environment is zoomed in, using the larger_animal dataset as an example. In the following, we provide a detailed explanation of the full methodology.
Overview. In each iteration $t\leq T$, the actor-critic agent generates a continuous vector “action” $\mathbf{a}$ (step 1). This action is then played on the environment, which projects $\mathbf{a}$ into the appropriate space using a fixed projection matrix $P$ to obtain $\mathbf{z}$. The environment then concatenates the projected vector $\mathbf{z}$ with a set of exemplars’ embeddings from $\mathcal{E}$ and feeds it into a white-box LLM $h$ (step 2). The white-box LLM produces a discrete prompt, $\tau$, which is evaluated using the validation dataset $\mathcal{V}$ based on the responses from the black-box LLM $f$ (step 3). The black-box LLM generates a prediction, which is then contrasted to the true labels of the validation examples, and a score function provides a final reward to the critic. This reward is used to compute the critic’s loss and update both the critic and actor networks accordingly.
[…]
By repeating steps 1 to 4 until the allowed budget $T$ is exhausted, the agent returns the best instruction prompt, denoted as $\tau^{\star}$. This prompt is then evaluated using the test set $\mathcal{T}$ on the black-box LLM, yielding the final score $\mathcal{S}^{\star}=\frac{1}{m^{\prime}}\sum_{i=1}^{m^{\prime}}q(f(\tau^{\star},x^{\prime}_{i}),y^{\prime}_{i})$ (using the evaluation template depicted in Fig. 1 bottom right).
Category Task Human Instruction ACING Instruction (Ours) Human Score Our Score
Spelling Second_word_letter Extract the second letter of the input word. Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: e 0.96 0.59
Syntax Negation Negate the input sentence. Flip the truth value of the statements in the input 0.81 0.82
Lexical Semantics Antonyms Write a word that means the opposite of the input word. Take a word and change it to its opposite 0.70 0.82
Synonyms Write a word with a similar meaning to the input word. Input a word that is a synonym for the word that was output 0.14 0.13
Phonetics Rhymes Write a word that rhymes with the input word. Input the word that the program thought I was inputting and then output the word that program thought I was inputting 0.61 1.00
Semantics Cause_and_effect Find which of the two given cause and effect sentences is the cause. Find the sentence that is the cause of the effect in the pair of sentences 0.97 0.92
Style Informal_to_formal Rephrase the sentence in formal language. Convert the input into output using the same word order and with the same meaning 0.63 0.5
Multi- lingual Translation_en-de Translate the word into German. Provide a translation for each word in the English text into German 0.81 0.84
Translation_en-es Translate the word into Spanish. Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.89 0.88
Translation_en-fr Translate the word into French. Create a program that would take an English word as input and output its French equivalent 0.86 0.87
GLUE Sentiment Determine whether a movie review is positive or negative. Classify each input as positive or negative based on the assessment of the corresponding movie 0.89 0.90
Sentence_similarity Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. Find a sentence pair that is probably not similar, and the output is 3 - probably 0.00 0.21
median score 0.81 0.83
average score 0.69 0.71
# best-performing tasks 5 7
Table 1: Tasks from the instruction-induction datasets where the human and ACING test scores differed. For each task, we provide the corresponding human instruction as proposed in Honovich et al. (2023) and our best discovered instruction. We tested these instructions on the test dataset and report their respective scores.
Category Task APE InstructZero INSTINCT ACING
Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00)
First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00)
Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.71 (0.06)
Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.74 (0.01)
Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.69 (0.31)
Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.84 (0.07)
Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.07)
Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.89 (0.01)
Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01)
Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.87 (0.02)
Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.83 (0.01)
Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.25 (0.00)
median score 0.49 0.66 0.64 0.71
# best-performing tasks 9 8 7 14
Table 2: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT. The bottom rows report the median score and total number of best-performing tasks for each method.
5 Experiments
We tackle instruction learning for ChatGPT. We conduct instruction induction tasks using 30 datasets from Honovich et al. (2023); Chen et al. (2024), along with a data summarization task from Gliwa et al. (2019). Furthermore, we compare the best-learned instructions from our ACING algorithm against human-provided instructions from Honovich et al. (2023), as shown in Table 1, with additional best-learned instructions in Table 9. Moreover, we compare ACING with three recent representative baselines for black-box instruction learning: APE (Zhou et al., 2023), InstructZero (Chen et al., 2024), and INSTINCT Lin et al. (2024), with results provided in Table 2.
Following previous works and to ensure fairness, we do not fine-tune the budget. Instead, like previous works, we consider a fixed budget of black-box LLM evaluations, set to $T=165$ Chen et al. (2024); Lin et al. (2024). Moreover, like previous works, we use Vicuna-13B Chiang et al. (2023) as the white-box LLM $h$. Furthermore, for previous works, we use their default (tuned) hyperparameters, including the intrinsic dimension $d^{\prime}=10$ and the number of soft tokens $N_{z}=5$. For fairness, we refrain from fine-tuning these parameters for our method and use the same values as in prior works. This ensures that our ACING algorithm searches in the same space, $[0,1]^{10}$, and uses the same total number of queries to the black-box LLM as APE, InstructZero, and INSTINCT for a fair comparison. For each algorithm, after identifying the best instruction using the validation set $\mathcal{V}$, we evaluate the discovered instruction on a separate test set $\mathcal{T}$ and report the test accuracy as the score.
We further conduct ablation studies on the key design choices for ACING, including the use of small learning budgets, providing reward plots for various selected tasks. Furthermore, we test the impact of using different white-box language models, intrinsic dimensions, and numbers of exemplars. More details on the experiments are provided in the Appendix.
5.1 ACING vs. Humans
We use human instructions provided in Honovich et al. (2023) for various instruction-induction tasks, test them, and compare their scores against those of our best instructions. The complete list of selected tasks is provided in Table 1 in the Appendix. Several instructions were solved by both human experts and our approach, i.e., achieving a perfect test score of 1.00. For clarity and brevity, we omit these results from the main paper and focus only on tasks where the scores of humans and our method differed, as shown in Table 1.
As shown in these tables, our approach outperforms human instructions in several tasks, sometimes by a significant margin. For example, in the rhyming task, our method achieved a 0.39 improvement in the test score. The human instruction for this task—Write a word that rhymes with the input word.—only resulted in a score of 0.61. In contrast, the best ACING instruction, which was an intriguing prompt: Input the word that the program thought I was inputting and then output the word that program thought I was inputting, achieved a perfect score of 1.00.
Furthermore, in the challenging sentence similarity task, while the human-provided instruction achieved a score of 0.00, our approach improved the test score by up to 0.21. However, the instruction yielding the highest score appeared to bias the prediction towards outputting “3” due to the following instruction: Find a sentence pair that is probably not similar, and the output is 3 - probably, which resulted in the best score. The second-best instruction, proposed by ACING, achieved an improvement of 0.14 over the human score and was as follows: Generate a probability score for each pair of sentences, where 1 - probably not means that the two sentences are not related, 0 - definitely not means that the two sentences are not related, and 3 - probably means that the two sentences are related. Overall, our approach recovers several human test scores while being fully automatic, and in many cases surpasses human performance in terms of the number of successful tasks, average score, and median score across tasks. For the curious reader, we provide all our best instructions along with their test scores in Table 10 in the Appendix.
5.2 ACING vs. Other Optimization Methods
We compare our method against recent baselines on the 30 instruction-induction datasets. Table 2 presents the average test accuracy (along with the standard deviation) over three independent runs, using three different seeds. For each seed, we selected the best instruction achieved by each method and evaluated it on the testing dataset. The table demonstrates that our method, ACING, outperforms others by achieving the highest accuracy in 14 out of 30 tasks, compared to INSTINCT, InstructZERO, and APE which succeeded in 9 tasks or less each. Additionally, ACING achieves the highest median accuracy across tasks, with a value of 0.71, which is  22 percentage points higher than APE. Further details, including the best prompt achieved for each task and corresponding test scores, are in the Appendix, Table 10. Moreover, we compare the performance of ACING against other methods on a summarization task using SAMSum dataset Gliwa et al. (2019) in Table 3, which shows that ACING outperforms the other methods.
In the following section, we demonstrate that the median score of ACING can be further improved by 5 percentage points using a simple two-phase exploration technique (splitting the budget), resulting in a 10-percentage-point improvement over the best baseline considered. Furthermore, in ablation studies, we show that fine-tuning the intrinsic dimensions $d^{{}^{\prime}}$ can lead to additional improvements.
Metric APE InstructZero INSTINCT ACING
ROUGE-1 0.35 (0.01) 0.33 (0.00) 0.36 (0.01) 0.37 (0.01)
ROUGE-2 0.12 (0.00) 0.11 (0.00) 0.14 (0.00) 0.14 (0.00)
ROUGE-L 0.25 (0.00) 0.24 (0.01) 0.27 (0.01) 0.28 (0.01)
Table 3: Average test performance (and standard deviations) for instruction optimization for summarization task using SAMSum dataset.
5.3 Splitting the Budget
Due to the stochastic nature of the black-box LLM, the same instruction may yield different rewards when evaluated by the LLM. For more robust prompt selection, the budget $T$ can be split into two parts: an actor-critic exploration phase (repeating steps 1 to 4), followed by a uniform exploration phase. In the latter phase, the top $p$ prompts discovered by the actor-critic are evaluated multiple times, $k$ times each, using the black-box LLM. Specifically, the first phase uses $T-p\cdot k$ API calls, while the second uses the remaining calls. The prompt with the highest average reward across repetitions can then be used at test time. In Table 5 in the Appendix, we show that ACING, using this budget split technique with an initial exploration budget of $150$ and a remaining budget of $15$ calls (allocated to uniformly explore the top $p=5$ prompts, each evaluated $k=3$ times), achieves improved median scores across tasks, resulting in 10 percentage points improvement over the best baseline considered and higher test accuracy on 15 tasks compared to other methods.
5.4 Ablation Studies
5.4.1 Using Smaller Budgets
In the main paper, we report the final test score after a fixed budget of 165 black-box LLM calls, mainly following previous work Lin et al. (2024); Chen et al. (2024), avoiding any potential advantage that could arise from optimizing this number. In the Appendix, we provide reward plots for the ACING approach, showing the best achieved reward over the calls. As shown in various plots in Figure 3, the ACING approach found the optimal instruction (achieving a reward of 1) within just a few black-box calls. Some tasks required fewer than 10 API calls to find the optimal instruction, as in the case of ‘active to passive’ and ‘letters list’, and fewer than 20 calls for tasks like ‘translation’ and ‘diff’. It can be seen that the vast majority of tasks achieved their best reward value within the first 60 to 80 calls, demonstrating that ACING can even be used for more constrained budgets.
5.4.2 Using a Different White-box Model
In the main paper, we present results using Vicuna as the white-box language model, following the setup of prior work. To evaluate the impact of different white-box models on the performance of our method, we also tested ACING with WizardLM-13B-v1.2 Xu et al. (2024) compared to Vicuna-13B-v1.3. The experiments in Table 8 indicate that WizardLM achieved a higher median test score across all tasks and excelled in a greater number of top-performing tasks, demonstrating that tuning the choice of the white-box model can improve results.
5.4.3 Using Different Action Dimensionalities
In the main paper, we present results using actions with dimension $d^{\prime}=10$, following the setup of prior work. To evaluate the performance of ACING with different dimensionalities, we conducted an experiment using $d^{\prime}\in\{5,10,20,40,100\}$, keeping other parameters fixed, for a budget of $165$ calls. The results, shown in Table 6, indicate that while the smallest dimension, $d^{\prime}=5$, recovered the best scores for some tasks, it generally has the lowest performance across most tasks. Furthermore, both $d^{\prime}=10$ and $d^{\prime}=20$ yield similar performance in terms of the number of best-performing tasks, indicating a low sensitivity to this parameter. For the larger dimension, $d^{\prime}=40$, the method achieved the highest number of best-performing tasks, showing that tuning (increasing) $d^{\prime}$ can improve results.
5.4.4 Using a Different Number of Exemplars
In the main paper, we present results using five exemplars, following the setup of prior work. To evaluate the performance of ACING with a smaller number of exemplars, we conducted an experiment using a single exemplar, i.e., $|\mathcal{E}|=1$. Results shown in Table 7 confirm that using five exemplars generally yield higher test scores and a greater number of best-performing tasks than a single exemplar. However, a single exemplar still performs well, matching the performance of five exemplars across several tasks (e.g., phonetics, summation, morpho-syntax, translation) and sometimes even outperforming them. Nonetheless, five exemplars offer an advantage, particularly for more cognitively complex tasks.
6 Related Work
Many early approaches such as AutoPrompt Shin et al. (2020), FluentPrompt Shi et al. (2023), and other soft prompt-based methods Lester et al. (2021); Li and Liang (2021); Zhong et al. (2021) focused on optimizing prompts for white-box LLMs and cannot be directly applied to black-box LLMs. Moreover, several techniques have also been proposed for optimizing prompts for grey-box LLMs: BBT Sun et al. (2022b), BBTv2 Sun et al. (2022a), and clip-tuning Chai et al. (2022) use evolutionary algorithms for this purpose. However, these methods still require access to input token embeddings and output logits, making them unsuitable for settings where only query access is allowed.
Other approaches Li and Liang (2021); Zhong et al. (2021) employ continuous soft prompts by fine-tuning parameters of specific input tokens, but these often generate human-unreadable prompts and are impractical for API-accessible LLMs. Some other methods Deng et al. (2022); Zhang et al. (2023), created for optimizing discrete prompts instead of continuous ones, can optimize discrete tokens but require access to the LLM’s output distribution, which is typically unavailable in black-box LLMs. Zhou et al. (2023) proposed generating candidate prompts through resampling without specific optimizations, while Guo et al. (2024) extended this model-free approach by using evolutionary algorithms. However, these methods usually require a large number of queries.
Instruction optimization in black-box LLMs recently focused on searching through the space of soft prompts, then relying on open-source white-box LLMs as a means to transform these soft prompts into discrete prompts. InstructZero Chen et al. (2024) learns discrete prompts for ChatGPT by searching through the continuous space of soft prompts using Bayesian optimization, which are then transformed into discrete prompts by a white-box LLM. INSTINCT Lin et al. (2024), a successor to this approach, relies on the same approach. However, instead of Bayesian optimization, it uses NeuralUCB Zhou et al. (2020), a contextual bandit (selection) algorithm that searches within a finite set of continuous actions generated using Sobol Eriksson et al. (2019). Our approach combines the hybrid strategy of using white-box and black-box LLMs with an actor-critic approach that operates in a stateless, infinite continuous-action space.
TRIPLE Shi et al. (2024) modeled the prompt learning problem as a best-arm-identification problem. However, it only considered the problem as a selection task, assuming a given pool of discrete prompts. In contrast, our method addresses both the generation and selection of prompts.
7 Conclusion
ACING presents a promising approach for learning prompts for black-box LLMs by framing it as a stateless continuous-action RL problem. Our method shows significant improvements over recent baseline methods. Evaluations across 31 datasets reveal that ACING enhances prompt quality, achieving a median score improvement of 10 percentage points and outperforming human-crafted instructions by up to 39 percentage points. This sets a new benchmark for automated instruction optimization in black-box LLMs.
8 Limitations
While our proposed results were achieved using default and commonly used values for hyperparameters in network design, learning rates, and following previous works for some common parameters, the results presented depend on these choices. These results could be further improved with fine-tuning, as demonstrated in the ablation studies (§ 5.4). Another limitation of this work, similar to previous studies, is the risk of overfitting when dealing with a limited number of validation examples. This can result in a prompt that minimizes the loss on the validation set but does not generalize well to the test dataset, leading to a discrepancy between training and test performance. Another limitation of this work, similar to previous work (Lin et al., 2024; Chen et al., 2024), is the reliance on a white-box model, where the choice of model may lead to performance variations, as demonstrated in § 5.4.2. Avoiding this limitation requires acting directly on the black-box LLM, which involves handling actions in a large discrete action space. This necessitates adapting current RL approaches for large discrete action spaces, as studied in Fourati et al. (2024) and the references therein.
Appendix A Experimental Details
A.1 Hyperparameters Remark
Across the diverse tasks, in the main paper, the same hyperparameters were used, which shows that the algorithm generalizes well across the 30 tasks without specifically tuning hyperparameters in each task. A summary of the key parameters cab be found in the following Table. An implementation of ACING is available at https://github.com/salmakh1/ACING.
Hyperparameter Choice
White-box $h$ Vicuna-13B
Actor-network $(1-1024-256-10)$
Critic-network $(10-128-128-1)$
Budget $T$ 165
Intrinsic (action) dimension $d^{\prime}$ 5
Number of soft tokens $N_{z}$ 5
Soft prompt dimension $d$ 5120 * $N_{z}$
Number of exemplars $|\mathcal{E}|$ 5
Number of tokens generated by Wb 64
Table 4: Key hyperparameters and their values.
A.2 Actor-Critic Details
Across all the tasks, we used three fully-connected layers for both the actor $(1-1024-256-10)$ and the critics $(10-128-128-1)$ networks, with learning rates fixed at $3\cdot 10^{-4}$ for each. We used two critic networks and consider their minimum as the actual critic. We learn the entropy parameter $\alpha$ using a learning rate of $9\cdot 10^{-4}$.
Appendix B ACING Rewards over the (Calls) Steps
In the main paper, we report the final test score after a fixed budget of 165 black-box LLM calls. In this section, we provide reward plots for the ACING approach, showing the best-achieved reward within the conducted calls. As shown in various plots in Figure 3, the ACING approach found the optimal prompt (achieving a reward of 1) within just a few black-box calls. Some tasks required fewer than 10 API calls to find the optimal instruction, such as for ‘active to passive’ and ‘letters list’, and fewer than 20 for tasks like ‘translation’ and ‘diff’. It can be seen that the vast majority of tasks achieved their best reward value within the first 60 to 80 calls, demonstrating that ACING can even be used for much more constrained budgets. The choice of 165 calls was mainly based on previous work Lin et al. (2024); Chen et al. (2024), avoiding any potential advantage that could come from optimizing this number.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Reward plots for running ACING on various selected tasks, showing the highest achieved reward on the y-axis until each API call (step), with the x-axis representing the number of API calls.
Appendix C ACING with Budget Splitting
Due to the stochastic nature of the black-box LLM, the same instruction may yield different rewards when evaluated by the LLM. To address this, we add a mechanism for more robust decision-making. The budget $T$ is split into two parts: an exploration phase where steps 1 to 4 are repeated, and an exploitation phase where the best $p$ prompts are evaluated multiple times, $k$ times each, using the black-box LLM. The exploration phase uses $T-p\cdot k$ API calls, with the remaining calls used for exploitation. Finally, the prompt with the highest average reward across repetitions is used at test time. In Table 5, we demonstrate that ACING, with an exploration budget of $T=150$ and the remaining 15 calls allocated to uniform exploration of the top $p=5$ prompts (evaluated $k=3$ times each), achieves improved median scores across tasks and higher test accuracy on 15 tasks compared to previous work, unlike without the splitting technique which achieved higher test accuracy on 14 tasks compared to the previous work.
Category Task APE InstructZero INSTINCT ACING150+15 ACING165
(budget splitting) (main paper)
Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00)
First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.40 (0.17) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.99 (0.01) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.82 (0.00) 0.71 (0.06)
Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.76 (0.06) 0.74 (0.01)
Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.12 (0.02) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.59 (0.05) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.60 (0.38) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.81 (0.30) 0.69 (0.31)
Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 0.97 (0.04) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.86 (0.06) 0.84 (0.07)
Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.03) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.76 (0.18) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.10 (0.01) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.48 (0.11) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.59 (0.05) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.54 (0.15) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.53 (0.34) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.27 (0.06) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.72 (0.02) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.08) 0.13 (0.07)
Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.88 (0.03) 0.89 (0.01)
Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01) 0.82 (0.01)
Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.86 (0.02) 0.87 (0.02)
Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.85 (0.02) 0.83 (0.01)
Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.29 (0.07) 0.25 (0.00)
median score 0.49 0.66 0.64 0.76 0.71
# best-performing tasks 7 8 8 15 14
Table 5: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT. The bottom rows report the median score and total number of best-performing tasks for each method.
Appendix D ACING with Different Intrinsic (Action) Dimensions
In the main paper, we present results using actions with a dimension of $d^{\prime}=10$, following the setup of prior work. To evaluate the performance of ACING across different dimensionalities, we conducted experiments with $d^{\prime}\in\{5,10,20,40\}$, keeping other parameters fixed, for a budget of $165$. We report the test results over different tasks and dimensionalities for a fixed seed. The results, shown in Table 6, indicate that while the smallest dimension, $d^{\prime}=5$, recovered the best scores for some tasks, it generally has the lowest performance across most tasks. Furthermore, both $d^{\prime}=10$ and $d^{\prime}=20$ yield similar performance in terms of the number of best-performing tasks (9-10 tasks), indicating low sensitivity to this parameter. For the much larger dimension, $d^{\prime}=40$, the method achieved the highest number of best-performing tasks (15 tasks), demonstrating improved performance with increased dimensionality. Further increasing the dimensionality to $d^{\prime}=100$ can still yield high results, outperforming $d^{\prime}\in{5,10,20}$. However, while it remarkably outperformed $d^{\prime}=40$ in some tasks, such as the second word letter task, synonyms, and antonyms, it only achieved 14 best-performing tasks overall, indicating similar but slightly lower performance than $d^{\prime}=40$.
Category Task $d^{{}^{\prime}}=5$ $d^{{}^{\prime}}=10$ $d^{{}^{\prime}}=20$ $d^{{}^{\prime}}=40$ $d^{{}^{\prime}}=100$
(main paper)
Spelling Letters_list 1.00 1.00 0.98 1.00 1.00
First_word_letter 1.00 1.00 0.97 1.00 1.00
Second_word_letter 0.23 0.91 0.30 0.29 0.92
Morpho-Syntax Singular_to_plural 0.99 0.99 1.00 1.00 1.00
Active_to_passive 1.00 1.00 1.00 1.00 1.00
Negation 0.82 0.80 0.84 0.81 0.70
Lexical Semantics Antonyms 0.73 0.76 0.76 0.82 0.84
Synonyms 0.12 0.13 0.14 0.14 0.34
Word_unscrambling 0.53 0.54 0.49 0.55 0.43
Phonetics Rhymes 0.95 0.36 0.94 1.00 1.00
Numerical Sum 0.99 0.79 1.00 0.99 1.00
Diff 0.89 1.00 1.00 1.00 1.00
Knowledge Larger_animal 0.79 0.94 0.93 0.65 0.68
Periodic_elements 1.00 0.98 0.94 0.98 0.98
Cognitive Tasks Cause_and_effect 0.64 0.52 0.92 0.56 0.56
Common_concept 0.12 0.23 0.11 0.12 0.02
Object_counting 0.51 0.39 0.48 0.59 0.44
Odd_one_out 0.60 0.64 0.64 0.68 0.26
Orthography_starts_with 0.11 0.65 0.59 0.61 0.71
Taxonomy_animal 0.79 0.68 0.59 0.85 0.97
Auto_categorization 0.30 0.28 0.13 0.33 0.32
Word_sorting 0.55 0.69 0.74 0.69 0.48
CLUE Sentence_similarity 0.00 0.21 0.00 0.14 0.07
Sentiment 0.91 0.88 0.86 0.91 0.80
Translation Num_to_verbal 0.99 1.00 1.00 1.00 1.00
Translation_en-de 0.83 0.81 0.81 0.80 0.81
Translation_en-es 0.89 0.90 0.91 0.91 0.86
Translation_en-fr 0.84 0.84 0.86 0.88 0.73
Style Informal_to_formal 0.54 0.40 0.51 0.49 0.50
Coding Auto_debugging 0.25 0.25 0.25 0.25 0.25
# best-performing tasks 8 9 10 15 14
Table 6: Average ACING test performance for a fixed random seed (0) with different soft prompt dimensions $d^{{}^{\prime}}$. The bottom row report the total number of best-performing tasks.
Appendix E ACING with Different Number of Exemplars
In this section, we test ACING with a single exemplar, in contrast to the main results in the paper, which use five exemplars for ACING and all other benchmarks. For these experiments, we fix all hyperparameters as in the main paper and run tests with a budget of $165$. Intuitively, providing more exemplars to the language model should facilitate prompt learning, so five exemplars are expected to yield better prompts than a single exemplar. Our experiments, summarized in Table 7, support this intuition. The results show that using five exemplars leads to higher test scores, as reflected in a greater number of best-performing tasks and an increase in median test scores across tasks. However, it is notable that performance did not decrease drastically with only one exemplar, suggesting that a single exemplar is sufficient to achieve decent results. In fact, across several tasks and categories (e.g., phonetics, summation, morpho-syntax, and translation), a single exemplar achieves the same performance of using five exemplars, and even outperforms the use of five exemplars in certain tasks. Nevertheless, using a single exemplar resulted in lower performance mainly in more cognitively challenging tasks, which is understandable, as more complex tasks are likely to benefit from additional exemplars.
Category Task ACING ($|\mathcal{E}|=1$) ACING ($|\mathcal{E}|=5$)
(main paper)
Spelling Letters_list 1.00 (0.00) 1.00 (0.00)
First_word_letter 0.99 (0.01) 1.00 (0.00)
Second_word_letter 0.19 (0.09) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 1.00 (0.00)
Negation 0.76 (0.08) 0.71 (0.06)
Lexical Semantics Antonyms 0.78 (0.05) 0.74 (0.01)
Synonyms 0.09 (0.03) 0.13 (0.02)
Word_unscrambling 0.41 (0.09) 0.50 (0.07)
Phonetics Rhymes 0.89 (0.08) 0.57 (0.31)
Numerical Sum 0.99 (0.01) 0.69 (0.31)
Diff 0.99 (0.01) 1.00 (0.00)
Knowledge Larger_animal 0.63 (0.17) 0.84 (0.07)
Periodic_elements 0.91 (0.08) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.51 (0.08) 0.69 (0.15)
Common_concept 0.16 (0.11) 0.19 (0.05)
Object_counting 0.26 (0.06) 0.41 (0.03)
Odd_one_out 0.64 (0.02) 0.64 (0.00)
Orthography_starts_with 0.06 (0.05) 0.60 (0.12)
Taxonomy_animal 0.63 (0.06) 0.71 (0.02)
Auto_categorization 0.01 (0.01) 0.29 (0.04)
Word_sorting 0.70 (0.02) 0.70 (0.03)
CLUE Sentence_similarity 0.07 (0.05) 0.13 (0.07)
Sentiment 0.70 (0.12) 0.89 (0.01)
Translation Num_to_verbal 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.72 (0.11) 0.82 (0.01)
Translation_en-es 0.88 (0.00) 0.87 (0.02)
Translation_en-fr 0.16 (0.04) 0.83 (0.01)
Style Informal_to_formal 0.42 (0.05) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00)
median score 0.67 0.71
# best-performing tasks 12 23
Table 7: Average ACING test performance (and standard deviations) across 3 random seeds comparing 1 exemplar versus 5 exemplars. The bottom rows report the median score and total number of best-performing tasks.
Appendix F ACING with Different White-box models
[…]
Category Task ACING (Vicuna) ACING (WizardLM)
(main paper)
Spelling Letters_list 1.00 (0.00) 1.00 (0.00)
First_word_letter 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.70 (0.15) 0.36 (0.18)
Morpho-Syntax Singular_to_plural 0.95 (0.03) 0.99 (0.00)
Active_to_passive 1.00 (0.00) 1.00 (0.00)
Negation 0.71 (0.06) 0.83 (0.00)
Lexical Semantics Antonyms 0.74 (0.01) 0.81 (0.02)
Synonyms 0.13 (0.02) 0.12 (0.03)
Word_unscrambling 0.50 (0.07) 0.57 (0.05)
Phonetics Rhymes 0.57 (0.31) 0.97 (0.04)
Numerical Sum 0.69 (0.31) 1.00 (0.00)
Diff 1.00 (0.00) 1.00 (0.00)
Knowledge Larger_animal 0.84 (0.07) 0.94 (0.01)
Periodic_elements 0.98 (0.00) 0.97 (0.02)
Cognitive Tasks Cause_and_effect 0.69 (0.15) 0.76 (0.20)
Common_concept 0.19 (0.05) 0.21 (0.05)
Object_counting 0.41 (0.03) 0.46 (0.07)
Odd_one_out 0.64 (0.00) 0.56 (0.11)
Orthography_starts_with 0.60 (0.12) 0.62 (0.03)
Taxonomy_animal 0.71 (0.02) 0.60 (0.32)
Auto_categorization 0.29 (0.04) 0.35 (0.03)
Word_sorting 0.70 (0.03) 0.61 (0.02)
CLUE Sentence_similarity 0.13 (0.07) 0.22 (0.04)
Sentiment 0.89 (0.01) 0.90 (0.02)
Translation Num_to_verbal 0.99 (0.01) 1.00 (0.00)
Translation_en-de 0.82 (0.01) 0.81 (0.01)
Translation_en-es 0.87 (0.02) 0.61 (0.38)
Translation_en-fr 0.83 (0.01) 0.83 (0.05)
Style Informal_to_formal 0.44 (0.05) 0.32 (0.19)
Coding Auto_debugging 0.25 (0.00) 0.38 (0.10)
median score 0.71 0.79
# best-performing tasks 14 21
Table 8: Average ACING test performance (and standard deviations) across 3 random seeds using Vicuna and WizardLM as white-box models. The bottom rows report the median score and total number of best-performing tasks.
Appendix G Demonstrations with Human Instructions
Category Task Demonstration Human Instruction Honovich et al. (2023) Human Score Our Score
Spelling First_word_letter cat $\rightarrow$ c Extract the first letter of the input word. 1.00 1.00
Second_word_letter cat $\rightarrow$ a Extract the second letter of the input word. 0.96 0.59
Letters_list cat $\rightarrow$ c a t Break the input word into letters, separated by spaces. 1.00 1.00
Morpho- syntax Singular_to_plural cat $\rightarrow$ cats Convert the input word to its plural form. 1.00 1.00
Active_to_passive The artist introduced the scientist. $\rightarrow$ The scientist was introduced by the artist. Write the input sentence in passive form. 1.00 1.00
Syntax Negation Time is finite $\rightarrow$ Time is not finite. Negate the input sentence. 0.81 0.82
Lexical Semantics Antonyms won $\rightarrow$ lost Write a word that means the opposite of the input word. 0.70 0.82
Synonyms alleged $\rightarrow$ supposed Write a word with a similar meaning to the input word. 0.14 0.13
Phonetics Rhymes sing $\rightarrow$ ring Write a word that rhymes with the input word. 0.61 1.00
Knowledge Larger_animal koala, snail $\rightarrow$ koala Write the larger of the two given animals. 0.94 0.94
Semantics Cause_and_effect Sentence 1: The soda went flat. Sentence 2: The bottle was left open. $\rightarrow$ The bottle was left open. Find which of the two given cause and effect sentences is the cause. 0.97 0.92
Common_concept guitars, pendulums, neutrinos $\rightarrow$ involve oscillations. Find a common characteristic for the given objects. 0.11 0.11
Style Informal_to_formal Please call once you get there $\rightarrow$ Please call upon your arrival. Rephrase the sentence in formal language. 0.63 0.5
Numerical Sum 22 10 $\rightarrow$ 32 Sum the two given numbers. 1.00 1.00
Diff 32 22 $\rightarrow$ 10 Subtract the second number from the first. 1.00 1.00
Num_to_Verbal 26 $\rightarrow$ twenty-six Write the number in English words. 1.00 1.00
Multi- lingual Translation_en-de game $\rightarrow$ Spiel Translate the word into German. 0.81 0.84
Translation_en-es game $\rightarrow$ juego Translate the word into Spanish. 0.89 0.88
Translation_en-fr game $\rightarrow$ jeu Translate the word into French. 0.86 0.87
GLUE Sentiment The film is small in scope, yet perfectly formed. $\rightarrow$ positive Determine whether a movie review is positive or negative. 0.89 0.90
Sentence_similarity Sentence 1: A man is smoking. Sentence 2: A man is skating. $\rightarrow$ 0 - definitely not Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. 0.00 0.21
median score 0.89 0.90
average score 0.78 0.79
# best-performing tasks 14 16
Table 9: Classified tasks into categories from the instruction-induction datasets. For each task, we provide a corresponding demonstration, with $\rightarrow$ separating the input from the output, along with its respective human instruction as proposed in Honovich et al. (2023). We tested these instructions, report their test scores, and compare them to our best test scores using ACING with Vicuna-13B as the white-box LLM.
Appendix H Our Best Learned Instructions
Task Best instruction Test Score
active_to_passive Change the input to match the output, but the output is already in the passive voice 1.00
antonyms Take a word and change it to its opposite 0.82
auto_categorization Match the input to the output, and the answer is:$Input$: Nature Nanotechnology, Annual Review of Biochemistry, and The Lancet Neurology $Output$: top journals $Input$: Jeans, Tops, and Suits $Output$: Apparel 0.34
auto_debugging Input the code into a Python interpreter and observe the output 0.375
cause_and_effect Find the sentence that is the cause of the effect in the pair of sentences 0.92
common_concept Make a connection between the input and output, but the connection is not clear 0.11
diff Find the difference between the two numbers 1.00
first_word_letter Create a function that takes a string as input and returns the first letter of the first word in the string 1.00
informal_to_formal Convert the input into output using the same word order and with the same meaning 0.50
larger_animal Create a program that takes two animals as input and outputs the animal that is bigger 0.94
letters_list Input the word "year" and the output was "y e a r" 1.00
negation Flip the truth value of the statements in the input 0.82
num_to_verbal Convert numbers to words 1.00
object_counting Provide a number that represents how many items are in the input 0.55
odd_one_out Find the word that does not belong in each group based on the given words 0.64
orthography_starts_with Find a word in the text that starts with the letter provided and to output that word 0.71
periodic_elements Find the name of the element based on its atomic number 1.00
rhymes Input the word that the program thought I was inputting and then output the word that program thought I was inputting 1.00
second_word_letter Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: 0.59
sentence_similarity Find a sentence pair that is probably not similar, and the output is 3 - probably 0.21
sentiment Classify each input as positive or negative based on the assessment of the corresponding movie 0.90
singular_to_plural Add the suffix -s to the end of the word to make it plural 1.00
sum Find the sum of the two numbers 1.00
synonyms Input a word that is a synonym for the word that was output 0.13
taxonomy_animal Make the AI generate a sequence of animals based on the input provided 0.75
translation_en-de Provide a translation for each word in the English text into German 0.84
translation_en-es Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.88
translation_en-fr Create a program that would take an English word as input and output its French equivalent 0.87
word_sorting Output the words in alphabetical order, but the output is not in alphabetical order 0.73
word_unscrambling Convert the input to a word that is a common English word 0.62
Table 10: The best instruction discovered by ACING for all the 30 instruction-induction tasks using with Vicuna-13B as the white-box LLM.

### v2

@@
-ACING: Actor-Critic for Instruction Learning in Black-Box Large Language Models
+ACING: Actor-Critic for Instruction Learning in Black-Box LLMs
 Abstract
-The effectiveness of Large Language Models (LLMs) in solving tasks vastly depends on the quality of the instructions, which often require fine-tuning through extensive human effort.
-This highlights the need for automated instruction optimization; however, this optimization is particularly challenging when dealing with black-box LLMs, where model parameters and gradients remain inaccessible.
-We propose ACING, a task-specific prompt optimization approach framed as a stateless continuous-action Reinforcement Learning (RL) problem, known as the continuum bandit setting.
-ACING leverages an actor-critic-based method to optimize prompts, learning from non-differentiable reward signals.
-We validate ACING by optimizing prompts for ChatGPT on 30 instruction-based tasks, as well as a summarization task.
-ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points compared to the best baseline considered.
-Furthermore, ACING not only recovers but also surpasses human-crafted expert instructions, achieving up to a 39 percentage point improvement over human benchmarks.
-ACING: Actor-Critic for Instruction Learning
-in Black-Box Large Language Models
-[…]
+The effectiveness of Large Language Models (LLMs) in solving tasks depends significantly on the quality of their instructions, which often require substantial human effort to craft.
+This underscores the need for automated instruction optimization.
+However, optimizing instructions is particularly challenging when working with black-box LLMs, where model parameters and gradients are inaccessible.
+We introduce ACING, an actor-critic reinforcement learning framework that formulates instruction optimization as a stateless, continuous-action problem, enabling exploration of infinite instruction spaces using only black-box feedback.
+ACING automatically discovers prompts that outperform human-written prompts in 76% of instruction-induction tasks, with gains of up to 33 points and a 10-point median improvement over the best automatic baseline in 33 tasks spanning instruction-induction, summarization, and chain-of-thought reasoning.
+Extensive ablations highlight its robustness and efficiency.
+An implementation of ACING is available at https://github.com/salmakh1/ACING.
+ACING: Actor-Critic for Instruction Learning in Black-Box LLMs
+Salma Kharrat          Fares Fourati KAUST {salma.kharrat, fares.fourati}@kaust.edu.sa          Marco Canini
 1 Introduction
-Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of tasks Zhao et al.
+Large Language Models (LLMs) have demonstrated impressive capabilities across tasks like classification, summarization, and reasoning Cobbe et al.
 (2021); Touvron et al.
+(2023); Zhao et al.
 (2024).
-This success is largely attributed to their strong instruction-following capabilities, which enable adaptation to diverse downstream applications Chen et al.
+A key driver of this success is their ability to follow natural language instructions, commonly known as prompts Chen et al.
+(2023a); Wei et al.
 (2022); Liu et al.
 (2023).
-These instructions, commonly referred to as prompts, play a crucial role in the performance of LLMs Wei et al.
-(2022); Zhu et al.
-(2024); Liu et al.
+Yet crafting effective prompts remains labor-intensive and brittle, particularly in black-box settings where model internals are inaccessible.
+A core challenge in prompt engineering lies in the extreme sensitivity of LLMs to subtle linguistic variations.
+Even minor changes in phrasing can lead to substantial performance shifts.
+As shown in Table˜1, a minor wording change yields a 12-point gain (using GPT-4o Hurst et al.
+(2024)) despite preserving semantics.
+Such prompt sensitivity is pervasive, underscoring the need for robust, automated prompt optimization methods.
+Recent work has explored automating prompt design to reduce human effort Reynolds and McDonell (2021); Mishra et al.
 (2021).
-Given the importance of prompts, researchers are increasingly interested in automating their optimization to reduce the need for manual adjustments, which is often a labor-intensive and costly process Reynolds and McDonell (2021); Mishra et al.
-(2021).
-As a result, developing efficient methods for automatic prompt optimization has become highly important to maximize LLMs performance.
-Several methods for automated prompt optimization have been proposed in the literature.
-Some focus on optimizing continuous prompts, known as soft prompts Zhong et al.
-(2021); Li and Liang (2021), which are typically fed into a LLM after its embedding layer since they do not represent meaningful language that the LLM can interpret directly.
-Other methods aim at optimizing discrete prompts, known as hard prompts, by either generating diverse prompts Zhou et al.
-(2023) or refining existing ones Pryzant et al.
-(2023).
-For both soft and hard prompts, some approaches rely on gradient-based methods Shin et al.
-(2020); Lester et al.
-(2021); Li and Liang (2021).
-However, these methods require access to the gradients of the LLMs and are therefore limited to white-box LLMs.
-Meanwhile, the most powerful LLMs today are typically black-box models (e.g., ChatGPT OpenAI (2023a) and GPT-4 OpenAI (2023b)).
-To overcome these challenges, recent studies have introduced gradient-free techniques for optimizing prompts in black-box LLMs Zhou et al.
-(2023); Prasad et al.
+Soft prompting techniques Li and Liang (2021); Lester et al.
+(2021) and heuristic-based discrete search methods Zhou et al.
 (2023); Pryzant et al.
-(2023).
-While promising, these approaches rely on heuristic local search methods and fail to fully utilize the observation history, including previously tested instructions and their outcomes, when selecting new instructions to test.
-Consequently, they are less effective at balancing the inherent exploration-exploitation trade-off, leading to inefficient querying.
-This inefficiency becomes problematic and impractical when API calls to black-box LLMs involve significant time and financial costs.
-A recent line of work has focused on optimizing instructions for black-box LLMs by searching through soft prompts and transforming them into discrete prompts using white-box LLMs.
-InstructZero Chen et al.
-(2024) optimizes discrete prompts for ChatGPT by applying Bayesian optimization over soft prompts.
-Building on the same line, INSTINCT Lin et al.
-(2024) replaces Bayesian optimization with NeuralUCB Zhou et al.
-(2020), a critic algorithm, to select actions, from a finite pool of soft prompts, initially generated with Sobol Eriksson et al.
-(2019).
-Inspired by the success of actor-critic methods for control problems Haarnoja et al.
-(2018, 2019) and generative adversarial networks Goodfellow et al.
-(2014), our approach combines the hybrid white-box/black-box framework with an actor-critic Reinforcement Learning (RL) algorithm, operating in a stateless, infinite continuous-action space, and constrained by a strict budget of API calls.
-The actor aims to maximize the critic’s value while remaining as stochastic as possible to enhance exploration and simultaneously exploit promising actions.
-This approach increases the likelihood of finding optimal prompts and enables more efficient use of the limited number of API calls.
-ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points.
-Furthermore, our optimized prompts not only recover but also surpass human-crafted expert instructions, achieving up to a 39 percentage point improvement against human benchmarks.
-An implementation of ACING is available at https://github.com/salmakh1/ACING.
+(2023) have shown promise, yet each faces key limitations in black-box settings.
+Soft prompts require access to the model internals, limiting them to white-box scenarios.
+Discrete search methods, in turn, often struggle to explore vast and nuanced instruction spaces efficiently.
+More recent hybrid methods combine white-box prompt generation with black-box evaluation Chen et al.
+(2023b); Lin et al.
+(2024b), but typically rely on finite candidate pools or rigid reward assumptions, constraining their applicability.
+To overcome these limitations, we introduce ACING, an actor-critic reinforcement learning (RL) framework for automated instruction optimization in black-box LLM settings.
+ACING formulates prompt optimization as a stateless, continuous-action RL problem within a continuum bandit environment.
+By learning a latent instruction space through an off-policy actor-critic algorithm, ACING efficiently explores infinite instruction candidates, using only black-box feedback for evaluation.
+A frozen white-box model is used as a decoder to convert latent vectors into discrete prompts, enabling both scalability and linguistic richness.
+The generated instructions are not only high-performing but also naturally interpretable, clear, and semantically aligned with the tasks, making them suitable for practical deployment.
+To effectively guide exploration under tight query budgets, ACING leverages entropy-regularized policy optimization, encouraging diversity in sampled prompts and improving the likelihood of discovering high-performing instructions.
+Source Instruction Score
+Human Write a word that means the opposite of the input word. 0.70
+ACING Take a word and change it to its opposite. 0.82
+Table 1: Our prompt vs human on GPT-4o for antonym task.
+Notably, to our knowledge, ACING is the first approach to apply off-policy, continuous-action actor-critic reinforcement learning to instruction learning in black-box LLMs, using lightweight actor and critic neural networks (rather than LLMs), making the approach both efficient and widely applicable.
+We evaluate ACING against state-of-the-art techniques, including Bayesian optimization, contextual bandits, and evolutionary strategies.
+Empirical results across 33 diverse tasks—including instruction induction, reasoning (including zero-shot Chain-of-Thought (CoT)), semantics, syntax, phonetics, translation, summarization, and code understanding—show that ACING outperforms both human-written prompts and strong automated baselines.
+Notably, it surpasses human instructions in 76% of instruction induction tasks, achieving gains of up to 33 points and a median improvement of 10 points over the best automatic baselines.
+In summary, our contributions are threefold:
+(1) An RL formulation of instruction learning: We propose a continuous-action actor-critic RL approach for instruction optimization in black-box LLMs, enabling scalable exploration of infinite instruction spaces using entropy-regularized policies.
+(2) Comprehensive validation.
+Across 33 diverse tasks, including instruction induction, zero-shot CoT reasoning, and summarization, ACING consistently outperforms both human-written and strong automated instructions, achieving statistically significant gains.
+(3) In-depth analysis and insights: Through extensive ablation studies, we analyze the impact of latent dimensionality, exemplar configuration, decoder architecture, and optimization budget, highlighting the robustness of the ACING approach.
+Additionally, human evaluation and automated readability analysis confirm the clarity and semantic faithfulness of the generated instructions.
+Figure 1: Pipeline of ACING.
+At each iteration, a soft prompt and task exemplars are fed to the white-box model to generate an instruction.
+This instruction queries the black-box LLM, whose outputs are scored.
+The resulting score is returned to the agent as a reward, which is used to update its networks and adjust its policy.
+Both LLMs remain frozen throughout.
 2 Problem Formulation
-2.1 Prompt Learning Objective
+2.1 Problem: Prompt Optimization for Black-Box LLMs
 We aim to improve the performance of a black-box LLM, denoted by $f$, which can only be accessed through its API, while its internal parameters remain unknown.
@@
 The black-box model $f$ processes an input formed by concatenating ($\oplus$) the prompt $\tau$ with the sentence $x$, producing a predicted output $\hat{y}=f(\tau\oplus x)$.
-More formally, the objective is to maximize the expected score of the LLM in solving the task $\mathcal{D}$, defined as:
+More formally, the objective is to maximize the expected score of the LLM in solving the task represented by the distribution $\mathcal{D}$, defined as:
 $\max_{\tau}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau\oplus x))\right].$ (1)
-To achieve this, we utilize a validation dataset $\mathcal{V}=\{(x_{j},y_{j})\}_{j=1}^{m}$, where each pair consists of an input sentence $x_{j}$ and its corresponding ground truth output $y_{j}$.
+We utilize a validation dataset $\mathcal{V}=\{(x_{j},y_{j})\}_{j=1}^{m}$, where each pair consists of an input sentence $x_{j}$ and its corresponding ground truth output $y_{j}$.
 Our objective is to find the prompt that maximizes the scoring function $q(\cdot,\cdot)$ across the validation dataset, where $q(\hat{y}_{j},y_{j})$ measures the quality of the predicted output $\hat{y}_{j}$ against the true output $y_{j}$.
@@
 The derived prompt $\tau^{\star}$ is then evaluated on a separate test set $\mathcal{T}=\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{m^{\prime}}$ to assess its generalization performance.
-2.2 Learning over a Continuous Space
+2.2 Reformulating Discrete Prompt Search as a Continuous Optimization Problem
 Directly optimizing the prompt $\tau$ for the black-box LLM model $f$ presents substantial challenges due to the discrete combinatorial nature of token selection in $\tau$.
-To mitigate this challenge, prior approaches like InstructZero Chen et al.
-(2024) and INSTINCT Lin et al.
-(2024) employ a white-box LLM, represented by $h$, to convert the discrete optimization problem into a continuous one by introducing a soft prompt vector $\mathbf{z}\in\mathbb{R}^{d}$, which is a continuous $d$-dimensional vector representing the token embedding of a set of virtual tokens.
-The white-box model $h$ has access to a dataset of exemplars, =$\mathcal{E}=\{(u_{j},v_{j})\}_{j=1}^{k}$, where each pair $(u_{j},v_{j})$ defines input-output text sequences that exemplify a downstream task.
-The white-box LLM $h$ serves as a proxy, mapping the soft prompt $\mathbf{z}$ into a discrete prompt $\tau$ suitable for input to the black-box LLM.
-Specifically, given the exemplars $\mathcal{E}$ and the vector $\mathbf{z}$, their concatenation is input to the white-box LLM, generating the discrete prompt $\tau(\mathbf{z})=h(\mathbf{z},\mathcal{E})$.
+To mitigate this challenge, similar to prior approaches Chen et al.
+(2024); Lin et al.
+(2024b); Hu et al.
+(2024) we employ a publicly available, open-source white-box model, represented by $h$, and introduce a soft prompt vector $\mathbf{z}\in\mathbb{R}^{d}$, which is a continuous $d$-dimensional vector representing the token embedding of a set of virtual tokens.
+The white-box model $h$, which remains entirely frozen, with no training or gradient updates, serves as a proxy mapping $\mathbf{z}$ into a discrete prompt $\tau$ for the black-box LLM.
+Given a dataset of exemplars, $\mathcal{E}=\{(u_{j},v_{j})\}_{j=1}^{k}$, where each pair $(u_{j},v_{j})$ defines input-output text sequences that exemplify a downstream task and the vector $\mathbf{z}$, their concatenation is input to the white-box model, generating the discrete prompt $\tau(\mathbf{z})=h(\mathbf{z},\mathcal{E})$.
 This generated prompt $\tau(\mathbf{z})$ is prepended to a test input $x_{j}$ from the validation set $\mathcal{V}$, and the combined input is provided to the black-box LLM $f$ to generate an output $\hat{y}_{j}=f(\tau(\mathbf{z})\oplus x_{j})$.
-The output $\hat{y}_{j}$ is then evaluated against the true output $y_{j}$ using the scoring function $q(\hat{y}_{j},y_{j})$.
-By using a fixed set of exemplars $\mathcal{E}$, the original discrete problem (Equation (1)) of finding the optimal prompt $\tau$ is effectively transformed into a continuous optimization problem over the soft prompt vector $\mathbf{z}$, as follows:
+The output $\hat{y}_{j}$ is then evaluated using the scoring function $q(\hat{y}_{j},y_{j})$.
+By using a fixed set of exemplars $\mathcal{E}$, the original discrete problem (Eq.
+(1)) of finding the optimal prompt $\tau$ is effectively transformed into a continuous optimization problem over the soft prompt vector $\mathbf{z}$, as follows:
 $\max_{\mathbf{z}\in\mathbb{R}^{d}}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau(\mathbf{z})\oplus x))\right].$ (2)
-2.3 Reducing the Continuous Space
-The soft prompt $\mathbf{z}$ is typically high-dimensional (e.g., $d=5120\times N_{z}$, with $N_{z}$ the number of virtual tokens, when $h$ is Vicuna-13B Chiang et al.
-(2023)), which makes it a challenging problem.
+The soft prompt $\mathbf{z}$ is typically high-dimensional.
 Therefore, we employ random projection techniques to reduce the input dimension as done in prior works Chen et al.
 (2024); Lin et al.
-(2024).
-Specifically, given a matrix $P\in\mathbb{R}^{d\times d^{\prime}}$ where $d^{\prime}\ll d$, with randomly sampled elements, and a continuous vector $\mathbf{a}\in\mathbb{R}^{d^{\prime}}$, the vector $\mathbf{z}=P\mathbf{a}$ is used as the soft prompt.
-By substituting $\mathbf{z}$ with $P\mathbf{a}$, the input variable to be optimized in Eq.
-2 becomes $\mathbf{a}$, thereby reducing the input dimension of the optimization problem to $d^{\prime}$:
+(2024b).
+Specifically, we sample a matrix $P\in\mathbb{R}^{d\times d^{\prime}}$ with entries from $\text{Uniform}(-1,1)$, and optimize a lower-dimensional vector $\mathbf{a}\in[0,1]^{d^{\prime}}$.
+The soft prompt is then given by $\mathbf{z}=P\mathbf{a}$, transforming the original problem into optimization over a compact and continuous space as follows:
 $\max_{\mathbf{a}\in\mathbb{R}^{d^{{}^{\prime}}}}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau(P\mathbf{a})\oplus x))\right].$ (3)
-For fairness with previous works, the reduced input dimension $d^{\prime}$, referred to as the intrinsic dimension, is chosen as $d^{\prime}=10$ Chen et al.
-(2024); Lin et al.
-(2024).
-An ablation study on the intrinsic dimension is discussed in § 5.4.3.
-Furthermore, to maintain consistency with these prior studies, we similarly use a projection matrix $P$ with elements uniformly and independently sampled from $\text{Uniform}(-1,1)$, and bound the vector such that $\mathbf{a}\in[0,1]^{d^{\prime}}$.
-3 Actor-Critic Approach for Instruction Learning in Black-Box LLMs
-3.1 Instruction Learning as an RL Problem
-In this work, we frame the prompt-learning problem as an agent navigating an infinite action space within a $d^{{}^{\prime}}$-dimensional unit hypercube, searching within an infinite number of actions $\mathbf{a}\in[0,1]^{d^{{}^{\prime}}}$, in contrast to previous works that consider a finite number of actions initially generated with Sobol Eriksson et al.
-(2019); Lin et al.
-(2024).
-As part of the environment, these continuous actions are projected into a higher-dimensional space (§ 2.3) and then transformed into discrete prompts, through a white-box LLM (§ 4).
-These discrete prompts are evaluated by a black-box LLM over a validation dataset, after which a reward is computed and returned to the agent.
-We consider a finite horizon $T$, representing the maximum number of actions played (i.e., the budget of evaluated instructions), after which the agent ceases exploration.
-The agent’s goal is to discover the optimal continuous action, $\mathbf{a}^{\star}$, which produces the optimal discrete prompt and the highest reward.
-Unlike in general RL settings, where the agent encounters multiple states with varying reward distributions, the agent we consider operates in a stateless, stochastic environment, commonly referred to as a stochastic bandit setting Slivkins (2019); Lattimore and Szepesvári (2020).
-In this context, at each iteration $t\leq T$ (where $t$ is incremented with every black-box LLM call), the agent’s action $\mathbf{a}_{t}$ generate stochastic rewards, $r_{t}\in[0,1]$, without affecting the reward distributions associated with those actions.
-The prompt-learning objective is modeled as a stochastic best-arm identification problem in a stateless, continuous action space RL, specifically in an infinite continuum bandit setting.
-Algorithm 1 Stateless Soft Actor-Critic
-0:  Evaluation budget $T$
-1:  Initialize $\boldsymbol{\theta}$, $\mathbf{w}$ (e.g., randomly)
-2:  for $t\leftarrow 1$ to $T$ do
-3:     Choose action $\mathbf{a}_{t}\sim\pi(\cdot\mid\boldsymbol{\theta})$
-4:     Take action $\mathbf{a}_{t}$ and observe reward $r_{t}$
-5:     $\mathbf{w}\leftarrow\mathbf{w}-\lambda\hat{\nabla}\mathbf{J}_{Q}(\mathbf{w})\quad$ (Equation (5))
-6:     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\beta\hat{\nabla}\mathbf{J}_{\pi}(\boldsymbol{\theta})\quad$ (Equation (6))
-7:     $\alpha\leftarrow\alpha-\gamma\hat{\nabla}\mathbf{J}_{\alpha}(\alpha)\quad$ (Equation (7))
-8:  end for
-3.2 Actor-Critic Continuum Bandits
-RL techniques are generally used to learn optimal policies and actions leading to best rewards.
-These techniques can be classified into value-based (critic-centric), policy-gradient (actor-centric), or hybrid approaches.
-Actor-critic algorithms merge the benefits of both by employing two key components: the actor, responsible for policy optimization, and the critic, which evaluates the policy (Sutton and Barto, 2018).
-This hybrid approach has demonstrated success in various domains, including continuous control problems Mnih et al.
+This optimization problem is central to our framework.
+The following section elaborates on our approach to solving it.
+For clarity, we summarize all symbols used in our actor-critic framework in Table 6.
+Figure 2: Illustration of the prompt generation and testing inside the environment using the larger_animal dataset as an example.
+3 Framework for Instruction Learning
+We formulate the problem of prompt learning for black-box LLMs as a RL problem, where the agent explores an infinite instruction space by sampling continuous actions $\mathbf{a}\in[0,1]^{d^{{}^{\prime}}}$.
+Each action corresponds to a latent prompt representation, which is mapped—via a fixed projection matrix and a decoder—into a discrete instruction.
+This instruction is then evaluated by the black-box LLM on a validation set to produce a reward indicating task performance.
+The stateless, stochastic setup places the problem in the continuum bandit regime, unlike traditional discrete multi-armed bandits Slivkins (2019); Lattimore and Szepesvári (2020).
+3.1 Actor-Critic for Instruction Optimization
+To efficiently learn in a high-dimensional and limited-feedback setting, we propose a stateless, off-policy actor-critic framework inspired by recent advances in continuous control Konda and Tsitsiklis (1999); Mnih et al.
 (2016); Haarnoja et al.
-(2018, 2019) and with generative adversarial networks (GANs) (Goodfellow et al., 2014), where a generator and a critic network collaboratively optimize performance.
-Figure 1: Pipeline of ACING.
-In each iteration, a soft prompt along with several examples of the target task are provided to the white-box LLM to generate an instruction.
-This instruction is then used to query the black-box LLM, which produces answers to the target task queries.
-The resulting score is returned to the agent as a reward, which is used to update its networks and adjust its policy.
-Both LLMs remain frozen throughout the process.
-Inspired by the success of actor-critic approaches, we introduce a (stateless) actor-critic algorithm, as provided in Algorithm 1, tailored to our infinite continuum bandit setting, enabling autonomous learning of effective prompts with a constrained evaluation budget and outperforming previous state-of-the-art black-box prompt learning approaches.
-Both the actor and the critic are parameterized by neural networks.
-In typical RL problems, the agent interacts with an environment with varying states, so its policy depends on the current state.
-That is, the policy network generates actions based on the current state $s$, $\pi(.|s;\theta)$, where $\theta$ denotes the parameters.
-However, in the problem we consider, the setting is stateless.
-Therefore, we define a policy network $\pi(.;\theta)$ with a constant input that produces a $d^{{}^{\prime}}$-dimensional vector as output $\mathbf{a}$, representing the continuous action.
-For the critic, we use a neural network with a $d^{{}^{\prime}}$-dimensional input (the action) and a single output that estimates the action’s quality, thereby assessing the quality of the policy that proposed it.
-The critic network’s parameters are trained to minimize the following:
+(2018, 2019).
+Our framework comprises lightweight networks: (1) a policy network (actor) $\pi(\cdot;\theta)$, which outputs latent action vectors corresponding to soft prompts, and (2) two critics to compute a value function $Q(\mathbf{a})=\min\{Q_{\mathbf{w_{1}}}(\mathbf{a}),Q_{\mathbf{w_{2}}}(\mathbf{a})\}$, which estimates the expected reward of a given action.
+Formally, the policy $\pi(\cdot;\theta)$ is modeled as a Gaussian distribution with a diagonal covariance matrix.
+The actor network outputs the mean and log standard deviation for each action dimension, and actions are sampled using the reparameterization trick.
+The critics, $Q_{\mathbf{w_{1}}}(.)$ and $Q_{\mathbf{w_{2}}}(.)$, are trained to minimize the mean squared error between predicted and observed rewards:
 $\min_{\mathbf{w}}J_{Q}(\mathbf{w})\triangleq\mathbb{E}_{\mathbf{a}\sim\mathcal{D}}\left[\frac{1}{2}\left(Q_{\mathbf{w}}\left(\mathbf{a}\right)-r\left(\mathbf{a}\right)\right)^{2}\right],$ (4)
 which can be optimized with stochastic gradients
-^$\hat{\nabla}_{\theta}J_{Q}(\mathbf{w})=\nabla_{\mathbf{w}}Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)\left(Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)-r\left(\mathbf{a}_{t}\right)\right),$ (5)
-where $Q$ is the critic-network with parameters $\mathbf{w}$. The actor-critic learning process alternates between policy-network and critic-network updates.
-3.3 Enhancing Exploration with Entropy
-In general, RL learning approaches aim to maximize the expected cumulative rewards up to a horizon $T$, expressed as $\max_{\pi(\cdot)}\sum_{t=1}^{T}\mathbb{E}_{a_{t}\sim\pi(\cdot)}\left[r(a_{t})\right].$ In contrast, black-box prompt learning, framed as a best-arm identification problem, seeks the best prompt without necessarily maximizing the expected cumulative rewards.
-To maximize the cumulative rewards, the agent must manage the exploration-exploitation trade-off.
-Conversely, for the best-arm identification problem, the agent should explore as extensively as possible within the $T$ evaluation budget to potentially identify the best action, leading to a pure exploration setting.
-Although it might seem that the agent should act purely at random to find the best action, some degree of exploitation may still be beneficial to guide the agent toward the optimal choice.
-To achieve a high exploration rate while exploiting prior data, we consider the maximum entropy objective, as discussed in previous RL works Ziebart (2010); Haarnoja et al.
+$\hat{\nabla}_{\mathbf{w}}J_{Q}(\mathbf{w})=\nabla_{\mathbf{w}}Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)\left(Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)-r\left(\mathbf{a}_{t}\right)\right).$ (5)
+Training alternates between actor and critic updates.
+Despite the lack of state transitions, this architecture proves effective: the critic generalizes reward signals in prompt space, stabilizes updates by reducing reward variance, and enhances robustness via a twin-critic setup to mitigate overestimation.
+Empirically, this actor-critic formulation consistently outperforms no-critic baselines.
+3.2 Enhancing Exploration via Entropy Regularization
+In our setting, the agent operates under a fixed evaluation budget of $T$ queries and must discover the best-performing instruction within this limit.
+This corresponds to a pure exploration regime, where broad coverage of the action space is essential.
+While random exploration may seem sufficient, effective learning often requires balancing exploration with exploitation of prior observations to guide the search toward high-reward regions.
+To encourage systematic exploration while leveraging past experience, we adopt the maximum entropy reinforcement learning framework Ziebart (2010); Haarnoja et al.
 (2018).
-This objective promotes stochastic policies by augmenting the reward with an entropy term, expressed as:
+This approach augments the expected reward with a policy entropy term, promoting stochasticity in the actor’s decisions.
+The objective becomes:
 $\min_{\theta}J_{\pi}(\theta)\triangleq\mathbb{E}_{\mathbf{a}\sim\pi(.;\theta)}\left[\alpha\log\left(\pi\left(\mathbf{a};\theta\right)\right)-Q_{\mathbf{w}}\left(\mathbf{a})\right)\right],$ (6)
-where the temperature $\alpha$ controls the stochasticity of the agent.
-The above objective is a simplification of the objective in the Soft Actor-Critic (SAC) algorithm Haarnoja et al.
-(2018), eliminating the state, as we consider a stateless environment and focus on instantaneous rather than cumulative rewards.
-The term $-\mathbb{E}_{a\sim\pi(.;\theta)}\left[\log\pi(a)\right]$ represents the entropy of the policy, which we seek to maximize alongside the critic values, as entropy maximization is crucial for encouraging the actor to explore different actions.
-A high $\alpha$ encourages exploration, whereas a lower value of $\alpha$ favors exploitation.
-Furthermore, to avoid fine-tuning, and to achieve adaptive exploration, the entropy coefficient $\alpha$ can be adjusted dynamically to maintain a target entropy level Haarnoja et al.
-(2019), $H_{\text{target}}$, ensuring sufficient exploration, as follows:
+where $\alpha$ is a temperature coefficient that governs the exploration-exploitation trade-off.
+A higher $\alpha$ encourages greater policy entropy, while a lower value biases the policy toward exploitation.
+This formulation simplifies the soft actor-critic objective Haarnoja et al.
+(2018), omitting state-related components and long-horizon returns, since our environment is stateless and rewards are immediate.
+The entropy term, $\mathbb{E}_{\mathbf{a}\sim\pi(\cdot;\theta)}[\log\pi(\mathbf{a})]$, incentivizes the actor to maintain a diverse action distribution, thereby avoiding premature convergence to suboptimal prompts.
+To avoid manual tuning of $\alpha$, we follow prior work Haarnoja et al. (2019) and adapt it to match a target entropy $H_{\text{target}}$, by minimizing:
 $\min_{\alpha}\mathbf{J}_{\alpha}\triangleq-\mathbb{E}_{a\sim\pi(.)}\left[\alpha\cdot\left(\log\pi(\mathbf{a};\theta)+H_{\text{target}}\right)\right].$ (7)
-We use a stochastic approximation of gradients, specifically the Adam algorithm Kingma and Ba (2017); Reddi et al.
-(2018), to minimize the losses via stochastic gradient descent.
-The learning rates $\lambda$, $\beta$, and $\gamma$ are used for the Q-network, the policy networks, and the temperature, respectively.
-Figure 2: Illustration of the prompt generation and testing inside the environment.
-4 Methodology
-The methodology of our proposed algorithm, ACING, is illustrated in Fig.
-1, showing the actor-critic interaction with the environment.
-In Fig.
-2, the environment is zoomed in, using the larger_animal dataset as an example.
-In the following, we provide a detailed explanation of the full methodology.
-Overview.
-In each iteration $t\leq T$, the actor-critic agent generates a continuous vector “action” $\mathbf{a}$ (step 1).
-This action is then played on the environment, which projects $\mathbf{a}$ into the appropriate space using a fixed projection matrix $P$ to obtain $\mathbf{z}$.
-The environment then concatenates the projected vector $\mathbf{z}$ with a set of exemplars’ embeddings from $\mathcal{E}$ and feeds it into a white-box LLM $h$ (step 2).
-The white-box LLM produces a discrete prompt, $\tau$, which is evaluated using the validation dataset $\mathcal{V}$ based on the responses from the black-box LLM $f$ (step 3).
-The black-box LLM generates a prediction, which is then contrasted to the true labels of the validation examples, and a score function provides a final reward to the critic.
-This reward is used to compute the critic’s loss and update both the critic and actor networks accordingly.
+We use stochastic gradient descent, specifically the Adam optimizer Kingma and Ba (2017); Reddi et al. (2018), to jointly update the policy, critic, and entropy temperature.
+3.3 Putting it All Together
+Fig. 1 illustrates our actor-critic framework and its interaction with the environment. Fig. 2 zooms in on an example from the larger_animal dataset.
+Overview: In each iteration $t\leq T$, the actor-critic agent generates a continuous vector “action” $\mathbf{a}$ (step 1).
+The action is then projected into the appropriate space using a fixed matrix $P$ to obtain $\mathbf{z}$.
+The environment then concatenates the projected vector $\mathbf{z}$ with a set of exemplars’ embeddings from $\mathcal{E}$ and feeds it into a white-box model $h$ (step 2).
+The white-box model produces a discrete prompt, $\tau$, which is evaluated using the validation dataset $\mathcal{V}$ based on the responses from the black-box LLM $f$ (step 3).
+The black-box LLM’s prediction is compared to true labels of the validation examples, and a score function provides a reward, used to update both the critic and actor networks accordingly.
 […]
@@
 […]
-By repeating steps 1 to 4 until the allowed budget $T$ is exhausted, the agent returns the best instruction prompt, denoted as $\tau^{\star}$.
-This prompt is then evaluated using the test set $\mathcal{T}$ on the black-box LLM, yielding the final score $\mathcal{S}^{\star}=\frac{1}{m^{\prime}}\sum_{i=1}^{m^{\prime}}q(f(\tau^{\star},x^{\prime}_{i}),y^{\prime}_{i})$ (using the evaluation template depicted in Fig.
-1 bottom right).
-Category Task Human Instruction ACING Instruction (Ours) Human Score Our Score
-Spelling Second_word_letter Extract the second letter of the input word.
-Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: e 0.96 0.59
-Syntax Negation Negate the input sentence. Flip the truth value of the statements in the input 0.81 0.82
-Lexical Semantics Antonyms Write a word that means the opposite of the input word. Take a word and change it to its opposite 0.70 0.82
-Synonyms Write a word with a similar meaning to the input word. Input a word that is a synonym for the word that was output 0.14 0.13
-Phonetics Rhymes Write a word that rhymes with the input word. Input the word that the program thought I was inputting and then output the word that program thought I was inputting 0.61 1.00
-Semantics Cause_and_effect Find which of the two given cause and effect sentences is the cause. Find the sentence that is the cause of the effect in the pair of sentences 0.97 0.92
-Style Informal_to_formal Rephrase the sentence in formal language. Convert the input into output using the same word order and with the same meaning 0.63 0.5
-Multi- lingual Translation_en-de Translate the word into German. Provide a translation for each word in the English text into German 0.81 0.84
-Translation_en-es Translate the word into Spanish. Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.89 0.88
-Translation_en-fr Translate the word into French. Create a program that would take an English word as input and output its French equivalent 0.86 0.87
-GLUE Sentiment Determine whether a movie review is positive or negative. Classify each input as positive or negative based on the assessment of the corresponding movie 0.89 0.90
+After $T$ iterations, the agent returns the best-performing prompt $\tau^{\star}$, which is evaluated on the test set $\mathcal{T}$ in the black-box LLM using the evaluation template shown in Fig.
+1 (bottom right).
+Category Task Human Instruction ACING Instruction (Ours) Human ACING
+Spelling Second_word_letter Extract the second letter of the input word. Input a word and output the letter that corresponds to the second letter in that word 0.96 (0.00) 0.92 (0.00)
+Syntax Negation Negate the input sentence. Flip the truth value of the statements in the input 0.81 (0.00) 0.82 (0.00)
+Lexical Semantics Antonyms Write a word that means the opposite of the input word. Take a word and change it to its opposite 0.70 (0.00) 0.83 (0.00)
+Synonyms Write a word with a similar meaning to the input word. Input a word that is a synonym for the word that was output 0.14 (0.01) 0.13 (0.00)
+Phonetics Rhymes Write a word that rhymes with the input word.
+Input the word that the program thought I was inputting and then output the word that program thought I was inputting 0.61 (0.01) 1.00 (0.00)
+Semantics Cause_and_effect Find which of the two given cause and effect sentences is the cause. Find the sentence that is the cause of the effect in the pair of sentences 0.97 (0.02) 0.90 (0.02)
+Style Informal_to_formal Rephrase the sentence in formal language. Convert the input into output using the same word order and with the same meaning 0.63 (0.00) 0.50 (0.00)
+Multi- lingual Translation_en-de Translate the word into German. Provide a translation for each word in the English text into German 0.81 (0.00) 0.84 (0.00)
+Translation_en-es Translate the word into Spanish. Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.89 (0.00) 0.88 (0.00)
+Translation_en-fr Translate the word into French. Create a program that would take an English word as input and output its French equivalent 0.86 (0.00) 0.87 (0.00)
+GLUE Sentiment Determine whether a movie review is positive or negative. Classify each input as positive or negative based on the assessment of the corresponding movie 0.89 (0.01) 0.91 (0.00)
 Sentence_similarity Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly.
-Find a sentence pair that is probably not similar, and the output is 3 - probably 0.00 0.21
-median score 0.81 0.83
-average score 0.69 0.71
+Find a sentence pair that is probably not similar, and the output is 3 - probably 0.00 (0.00) 0.21 (0.00)
+median score 0.81 0.86
 # best-performing tasks 5 7
@@
 (2023) and our best discovered instruction.
-We tested these instructions on the test dataset and report their respective scores.
-Category Task APE InstructZero INSTINCT ACING
-Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00)
-First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
-Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.70 (0.15)
-Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.95 (0.03)
-Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00)
-Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.71 (0.06)
-Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.74 (0.01)
-Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.13 (0.02)
-Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.50 (0.07)
-Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.57 (0.31)
-Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.69 (0.31)
-Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 1.00 (0.00)
-Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.84 (0.07)
-Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.00)
-Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.69 (0.15)
-Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.19 (0.05)
-Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.41 (0.03)
-Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.64 (0.00)
-Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.60 (0.12)
-Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.71 (0.02)
-Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.29 (0.04)
-Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.70 (0.03)
-CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.07)
-Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.89 (0.01)
-Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 0.99 (0.01)
-Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01)
-Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.87 (0.02)
-Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.83 (0.01)
-Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05)
-Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.25 (0.00)
-median score 0.49 0.66 0.64 0.71
-# best-performing tasks 9 8 7 14
-Table 2: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT.
-The bottom rows report the median score and total number of best-performing tasks for each method.
+We tested these instructions on the test dataset and report the average score (with standard deviation) over 3 repetitions.
+Category Task APE EvoPrompt InstructZero INSTINCT ACING
+Spelling Letters_list 0.59 (0.02) 0.97 (0.03) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00)
+First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
+Second_word_letter 0.00 (0.00) 0.63 (0.17) 0.35 (0.09) 0.39 (0.28) 0.70 (0.15)
+Morpho-Syntax Negation 0.79 (0.00) 0.84 (0.02) 0.65 (0.10) 0.58 (0.22) 0.71 (0.06)
+Lexical Semantics Synonyms 0.14 (0.01) 0.19 (0.07) 0.22 (0.11) 0.19 (0.08) 0.13 (0.02)
+Word_unscrambling 0.54 (0.00) 0.44 (0.06) 0.59 (0.06) 0.54 (0.02) 0.50 (0.07)
+Phonetics Rhymes 0.59 (0.01) 0.52 (0.05) 0.99 (0.01) 0.36 (0.04) 0.57 (0.31)
+Numerical Sum 0.87 (0.01) 1.00 (0.00) 1.00 (0.00) 0.70 (0.21) 1.00 (0.00)
+Diff 0.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.93 (0.09) 1.00 (0.00)
+Knowledge Larger_animal 0.72 (0.02) 0.58 (0.06) 0.63 (0.07) 0.81 (0.09) 0.84 (0.07)
+Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.48 (0.10) 0.52 (0.09) 0.55 (0.11) 0.69 (0.15)
+Common_concept 0.03 (0.02) 0.17 (0.00) 0.14 (0.04) 0.09 (0.04) 0.19 (0.05)
+Object_counting 0.30 (0.02) 0.50 (0.06) 0.38 (0.06) 0.40 (0.12) 0.41 (0.03)
+Odd_one_out 0.32 (0.02) 0.64 (0.04) 0.57 (0.02) 0.25 (0.18) 0.64 (0.00)
+Orthography_starts_with 0.23 (0.01) 0.47 (0.02) 0.41 (0.09) 0.54 (0.06) 0.60 (0.12)
+Taxonomy_animal 0.02 (0.02) 0.38 (0.15) 0.67 (0.14) 0.85 (0.06) 0.71 (0.02)
+Auto_categorization 0.31 (0.01) 0.20 (0.03) 0.29 (0.02) 0.07 (0.07) 0.29 (0.04)
+Word_sorting 0.58 (0.01) 0.01 (0.00) 0.64 (0.05) 0.23 (0.20) 0.70 (0.03)
+CLUE Sentence_similarity 0.00 (0.00) 0.05 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.07)
+Translation Num_to_verbal 0.13 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.99 (0.01)
+Translation_en-es 0.86 (0.01) 0.76 (0.00) 0.67 (0.24) 0.89 (0.00) 0.87 (0.02)
+Style Informal_to_formal 0.57 (0.01) 0.50 (0.02) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05)
+Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.25 (0.00)
+median score 0.31 0.50 0.59 0.54 0.69
+# best-performing tasks 3 7 8 4 13
+Table 3: Average test performance (with standard deviations) over 3 seeds comparing ACING to APE (Zhou et al., 2023), EvoPrompt (Guo et al., 2024), InstructZero (Chen et al., 2024), and INSTINCT (Lin et al., 2024b) on the 23 most challenging tasks (where at least one method has score < 0.7).
+Bottom rows show median scores and the number of best-performing tasks.
 4 Experiments
-We tackle instruction learning for ChatGPT.
-We conduct instruction induction tasks using 30 datasets from Honovich et al.
+We focus on instruction learning for ChatGPT OpenAI (2023b), with additional analysis on GPT-4 OpenAI (2023c) and GPT-4o Hurst et al.
+(2024) as representative black-box LLMs.
+We conduct instruction induction tasks using 30 datasets spanning several diverse categories from Honovich et al.
 (2023); Chen et al.
-(2024), along with a data summarization task from Gliwa et al.
+(2024), zero-shot CoT reasoning datasets (GSM8K Cobbe et al.
+(2021), AQUARAT Ling et al.
+(2017)), and summarization on SAMSum Gliwa et al.
 (2019).
-Furthermore, we compare the best-learned instructions from our ACING algorithm against human-provided instructions from Honovich et al.
-(2023), as shown in Table 1, with additional best-learned instructions in Table 9.
-Moreover, we compare ACING with three recent representative baselines for black-box instruction learning: APE (Zhou et al., 2023), InstructZero (Chen et al., 2024), and INSTINCT Lin et al.
-(2024), with results provided in Table 2.
-Following previous works and to ensure fairness, we do not fine-tune the budget.
-Instead, like previous works, we consider a fixed budget of black-box LLM evaluations, set to $T=165$ Chen et al.
-(2024); Lin et al.
+In §4.1, we compare ACING’s best-learned instructions against human-written prompts from Honovich et al.
 (2023).
-Moreover, like previous works, we use Vicuna-13B Chiang et al.
-(2023) as the white-box LLM $h$.
-Furthermore, for previous works, we use their default (tuned) hyperparameters, including the intrinsic dimension $d^{\prime}=10$ and the number of soft tokens $N_{z}=5$.
-For fairness, we refrain from fine-tuning these parameters for our method and use the same values as in prior works.
-This ensures that our ACING algorithm searches in the same space, $[0,1]^{10}$, and uses the same total number of queries to the black-box LLM as APE, InstructZero, and INSTINCT for a fair comparison.
-For each algorithm, after identifying the best instruction using the validation set $\mathcal{V}$, we evaluate the discovered instruction on a separate test set $\mathcal{T}$ and report the test accuracy as the score.
-We further conduct ablation studies on the key design choices for ACING, including the use of small learning budgets, providing reward plots for various selected tasks.
-Furthermore, we test the impact of using different white-box language models, intrinsic dimensions, and numbers of exemplars.
-More details on the experiments are provided in the Appendix.
+We also benchmark, in §4.2, against four recent black-box instruction optimization methods: APE Zhou et al.
+(2023), EvoPrompt Guo et al.
+(2024), InstructZero Chen et al.
+(2024), and INSTINCT Lin et al.
+(2024b).
+Furthermore, we study the interpretability and clarity of generated instructions (§4.3) and conduct ablation studies (§4.4) to examine the impact of the different design choices.
+To ensure fairness and comparability with other automatic approaches, we adopt a fixed evaluation budget of black-box queries T=165T=165, [^1] See Appendix F.3 for reward plots where ACING often peaks well before the budget is exhausted.
+consistent with prior work.
+Moroever, ACING uses an off-the-shelf, publicly available white-box decoder to map latent vectors to prompts.
+Moreover, all learning and evaluation occur solely through black-box interactions, which is consistent with other leading methods (e.g., INSTINCT, InstructZero).
+To ensure fairness, we fix the same decoder Vicuna-13B Chiang et al.
+(2023), as used in their analyses across all methods, while in ablations we compare with WizardLM-13B Xu et al.
+(2024).
+The experimental setup isolates ACING’s core contribution, ensuring that performance gains are not attributable to decoder selection or tuning.
+Further details, including hyperparameters, are provided in Appendix B.
 4.1 ACING vs. Humans
-We use human instructions provided in Honovich et al.
-(2023) for various instruction-induction tasks, test them, and compare their scores against those of our best instructions.
-The complete list of selected tasks is provided in Table 1 in the Appendix.
-Several instructions were solved by both human experts and our approach, i.e., achieving a perfect test score of 1.00.
-For clarity and brevity, we omit these results from the main paper and focus only on tasks where the scores of humans and our method differed, as shown in Table 1.
-As shown in these tables, our approach outperforms human instructions in several tasks, sometimes by a significant margin.
-For example, in the rhyming task, our method achieved a 0.39 improvement in the test score.
-The human instruction for this task—Write a word that rhymes with the input word.—only resulted in a score of 0.61.
-In contrast, the best ACING instruction, which was an intriguing prompt: Input the word that the program thought I was inputting and then output the word that program thought I was inputting, achieved a perfect score of 1.00.
-Furthermore, in the challenging sentence similarity task, while the human-provided instruction achieved a score of 0.00, our approach improved the test score by up to 0.21.
-However, the instruction yielding the highest score appeared to bias the prediction towards outputting “3” due to the following instruction: Find a sentence pair that is probably not similar, and the output is 3 - probably, which resulted in the best score.
-The second-best instruction, proposed by ACING, achieved an improvement of 0.14 over the human score and was as follows: Generate a probability score for each pair of sentences, where 1 - probably not means that the two sentences are not related, 0 - definitely not means that the two sentences are not related, and 3 - probably means that the two sentences are related.
-Overall, our approach recovers several human test scores while being fully automatic, and in many cases surpasses human performance in terms of the number of successful tasks, average score, and median score across tasks.
-For the curious reader, we provide all our best instructions along with their test scores in Table 10 in the Appendix.
+We compare instructions found by ACING against human-authored prompts from Honovich et al.
+(2023) across a broad range of instruction-induction tasks.
+Table 2 highlights only those tasks where test performance differed between the two.
+The full set of tasks is included in Appendix G.
+Acing not only matches but often surpasses human-written instructions—often by substantial margins.
+For instance, in the Antonyms task, the human instruction (“Write a word that means the opposite of the input word”) scores 0.70.
+ACING improves this to 0.82 with a more direct and actionable phrasing: “Take a word and change it to its opposite.” The formulation is simpler yet effective.
+Consider the rhyming task, where the human instruction—“Write a word that rhymes with the input word”—yields a score of 0.61.
+ACING significantly improves performance, achieving a perfect score of 1.00 with an alternative phrasing.
+This highlights ACING’s ability to discover high-performing instructions that align closely with the underlying model behavior.
+In the sentence similarity task, where the human instruction results in a score of 0.00, ACING raises performance to 0.21.
+The highest-scoring ACING instruction introduces a mild incline toward a mid-scale output (“3 - probably”), but its phrasing remains semantically valid and interpretable.
+A second-best instruction, unbiased, still improves upon the human-written version by 0.14 points.
+On average, Acing improves the median task score from 0.81 to 0.86 and outperforms human-written instructions on 7 out of 12 tasks in this subset—underscoring its potential as a practical and effective alternative to manual prompt engineering.
+Further generated instructions are in Appendix I.
+Furthermore, extended results with GPT-4o are provided in Table˜11 in Appendix˜E.
 4.2 ACING vs. Other Optimization Methods
-We compare our method against recent baselines on the 30 instruction-induction datasets.
-Table 2 presents the average test accuracy (along with the standard deviation) over three independent runs, using three different seeds.
+Instruction-induction datasets: In Table 3, we show tasks from Honovich et al.
+(2023) where at least one method failed to achieve 70% score (full results on all 30 tasks can be found in Appendix˜C).
+The table shows the average test accuracy (along with the standard deviation) over three independent runs, using three different seeds.
 For each seed, we selected the best instruction achieved by each method and evaluated it on the testing dataset.
-The table demonstrates that our method, ACING, outperforms others by achieving the highest accuracy in 14 out of 30 tasks, compared to INSTINCT, InstructZERO, and APE which succeeded in 9 tasks or less each.
-Additionally, ACING achieves the highest median accuracy across tasks, with a value of 0.71, which is  22 percentage points higher than APE.
-Further details, including the best prompt achieved for each task and corresponding test scores, are in the Appendix, Table 10.
-Moreover, we compare the performance of ACING against other methods on a summarization task using SAMSum dataset Gliwa et al.
-(2019) in Table 3, which shows that ACING outperforms the other methods.
-In the following section, we demonstrate that the median score of ACING can be further improved by 5 percentage points using a simple two-phase exploration technique (splitting the budget), resulting in a 10-percentage-point improvement over the best baseline considered.
-Furthermore, in ablation studies, we show that fine-tuning the intrinsic dimensions $d^{{}^{\prime}}$ can lead to additional improvements.
-Metric APE InstructZero INSTINCT ACING
-ROUGE-1 0.35 (0.01) 0.33 (0.00) 0.36 (0.01) 0.37 (0.01)
-ROUGE-2 0.12 (0.00) 0.11 (0.00) 0.14 (0.00) 0.14 (0.00)
-ROUGE-L 0.25 (0.00) 0.24 (0.01) 0.27 (0.01) 0.28 (0.01)
-Table 3: Average test performance (and standard deviations) for instruction optimization for summarization task using SAMSum dataset.
-5.3 Splitting the Budget
-Due to the stochastic nature of the black-box LLM, the same instruction may yield different rewards when evaluated by the LLM.
-For more robust prompt selection, the budget $T$ can be split into two parts: an actor-critic exploration phase (repeating steps 1 to 4), followed by a uniform exploration phase.
-In the latter phase, the top $p$ prompts discovered by the actor-critic are evaluated multiple times, $k$ times each, using the black-box LLM.
-Specifically, the first phase uses $T-p\cdot k$ API calls, while the second uses the remaining calls.
-The prompt with the highest average reward across repetitions can then be used at test time.
-In Table 5 in the Appendix, we show that ACING, using this budget split technique with an initial exploration budget of $150$ and a remaining budget of $15$ calls (allocated to uniformly explore the top $p=5$ prompts, each evaluated $k=3$ times), achieves improved median scores across tasks, resulting in 10 percentage points improvement over the best baseline considered and higher test accuracy on 15 tasks compared to other methods.
+The results demonstrate that our method, ACING, outperforms the others, achieving the highest accuracy in 13 out of the remaining 23 tasks, compared to INSTINCT, InstructZERO, EvoPrompt, and APE, which succeeded in 8 tasks or fewer.
+Additionally, ACING achieves the highest median accuracy across tasks, with a value of 0.69, which is approximately 10 percentage points higher than the best baseline.
+The score types can be found in Table 8.
+Moreover, the best prompt achieved for each task and the corresponding test scores, can be found in Table 21 in Appendix˜C.
+Further analyses on GPT-4 can be found in the Appendix.
+D.
+CoT datasets.
+We evaluate our method on two zero-shot CoT reasoning datasets: GSM8K Cobbe et al.
+(2021) and AQUA-RAT Ling et al.
+(2017).
+Prior work Kojima et al.
+(2022) shows that simple CoT prompts can enhance LLM performance.
+Using 100 steps (and other settings as above), ACING achieves the best results on both, demonstrating strong CoT reasoning capability.
+Summarization dataset.
+We compare the performance of ACING with other methods on summarization tasks using the SAMSum dataset Gliwa et al.
+(2019).
+The results, presented in Table 4, show that ACING outperforms the other methods across the three metrics considered: ROUGE-1, ROUGE-2, and ROUGE-L Lin (2004).
+Statistical significance test.
+We conduct a Wilcoxon signed-rank test Wilcoxon (1992), a non-parametric test.
+The results confirm that ACING significantly outperforms all baselines across tasks, with p = 0.0005 (APE), 0.0041 (INSTINCT), 0.0092 (EvoPrompt), and 0.0335 (InstructZero), all below the standard 0.05 threshold, indicating that the observed gains are statistically significant.
+Metric APE EvoPompt InstructZero INSTINCT ACING
+ROUGE-1 0.35 (0.01) 0.35 (0.01) 0.33 (0.00) 0.36 (0.01) 0.37 (0.01)
+ROUGE-2 0.12 (0.00) 0.12 (0.00) 0.11 (0.00) 0.14 (0.00) 0.14 (0.00)
+ROUGE-L 0.25 (0.00) 0.26 (0.00) 0.24 (0.01) 0.27 (0.01) 0.28 (0.01)
+Table 4: Average test performance (and standard deviations) for summarization task using SAMSum dataset.
+Results breakdown.
+To better understand ACING’s strengths, we grouped the tasks into various categories.
+On cognitive tasks, requiring complex reasoning, ACING demonstrates the clearest advantage.
+It ranks first on 5 out of 8 such tasks from Table 3, including Cause_and_effect (0.69, +14%), Word_sorting (0.70, +6%), Odd_one_out (0.64, tied), and Orthography_starts_with (0.60, +6%).
+As shown in Fig.
+3 (ordered left to right by median ranking), ACING achieves the best median rank and shows a strong skew toward top rankings in this category.
+Furthermore, ACING shows the best performance in the zero-shot CoT reasoning tasks (Table˜5).
+In symbolic manipulation, ACING again performs best, achieving perfect scores on Letters_list and First_word_letter, and leading on Second_word_letter (0.70 vs.
+0.63).
+This highlights its precision in low-level, structured tasks.
+ACING is competitive but not dominant in tasks involving lexical semantics, world knowledge, and translation—e.g., Informal_to_formal (0.44 vs.
+0.57), and Translation_en-es (0.87 vs.
+0.89).
+Finally, it consistently outperforms the other methods in summarization.
+Method Dataset Best Zero-Shot Instruction Score
+Kojima et al. (2022) GSM8K Let’s think step by step. 0.72
+INSTINCT Lin et al. (2024b) GSM8K Let’s use our creativity to find the solution. 0.75
+ACING (Ours) GSM8K Let’s use our math skills to conquer this challenge. 0.76
+Kojima et al. (2022) AQUA-RAT Let’s think step by step. 0.59
+INSTINCT Lin et al. (2024b) AQUA-RAT Let’s break it down. 0.59
+ACING (Ours) AQUA-RAT Let’s use the power of substitution to solve this problem. 0.63
+Table 5: Best zero-shot instructions and corresponding scores for different methods and datasets.
+4.3 Instruction Clarity and Readability
+To evaluate the interpretability and clarity of generated instructions, we assess whether ACING-generated prompts are understandable and task-aligned.
+We conduct a human evaluation with 26 participants who rated the clarity and alignment of generated instructions from Table˜2 on a 5-point Likert scale.
+The results show that participants found the prompts generally clear, with a median score of 3.9/5 (see Appendix H for the protocol and statistics).
+In parallel, we use standard readability metrics to automatically assess the generated prompts.
+The results show a median Flesch Reading Ease (FRE) of 70.8, Flesch-Kincaid Grade Level (FKG) of 7.0, and Coleman-Liau Index (CLI) of 7.3.
+These scores correspond to mid-grade readability (7th–8th grade), indicating that the generated instructions are accessible to a broad user base.
 4.4 Ablation Studies
-5.4.1 Using Smaller Budgets
-In the main paper, we report the final test score after a fixed budget of 165 black-box LLM calls, mainly following previous work Lin et al.
-(2024); Chen et al.
-(2024), avoiding any potential advantage that could arise from optimizing this number.
-In the Appendix, we provide reward plots for the ACING approach, showing the best achieved reward over the calls.
-As shown in various plots in Figure 3, the ACING approach found the optimal instruction (achieving a reward of 1) within just a few black-box calls.
-Some tasks required fewer than 10 API calls to find the optimal instruction, as in the case of ‘active to passive’ and ‘letters list’, and fewer than 20 calls for tasks like ‘translation’ and ‘diff’.
-It can be seen that the vast majority of tasks achieved their best reward value within the first 60 to 80 calls, demonstrating that ACING can even be used for more constrained budgets.
-5.4.2 Using a Different White-box Model
-In the main paper, we present results using Vicuna as the white-box language model, following the setup of prior work.
-To evaluate the impact of different white-box models on the performance of our method, we also tested ACING with WizardLM-13B-v1.2 Xu et al.
-(2024) compared to Vicuna-13B-v1.3.
-The experiments in Table 8 indicate that WizardLM achieved a higher median test score across all tasks and excelled in a greater number of top-performing tasks, demonstrating that tuning the choice of the white-box model can improve results.
-5.4.3 Using Different Action Dimensionalities
-In the main paper, we present results using actions with dimension $d^{\prime}=10$, following the setup of prior work.
-To evaluate the performance of ACING with different dimensionalities, we conducted an experiment using $d^{\prime}\in\{5,10,20,40,100\}$, keeping other parameters fixed, for a budget of $165$ calls.
-The results, shown in Table 6, indicate that while the smallest dimension, $d^{\prime}=5$, recovered the best scores for some tasks, it generally has the lowest performance across most tasks.
-Furthermore, both $d^{\prime}=10$ and $d^{\prime}=20$ yield similar performance in terms of the number of best-performing tasks, indicating a low sensitivity to this parameter.
-For the larger dimension, $d^{\prime}=40$, the method achieved the highest number of best-performing tasks, showing that tuning (increasing) $d^{\prime}$ can improve results.
-5.4.4 Using a Different Number of Exemplars
-In the main paper, we present results using five exemplars, following the setup of prior work.
-To evaluate the performance of ACING with a smaller number of exemplars, we conducted an experiment using a single exemplar, i.e., $|\mathcal{E}|=1$.
-Results shown in Table 7 confirm that using five exemplars generally yield higher test scores and a greater number of best-performing tasks than a single exemplar.
-However, a single exemplar still performs well, matching the performance of five exemplars across several tasks (e.g., phonetics, summation, morpho-syntax, translation) and sometimes even outperforming them.
-Nonetheless, five exemplars offer an advantage, particularly for more cognitively complex tasks.
+We present detailed ablation studies on key design choices, summarize some below, and provide full results and plots in Appendices F.1F.7.
+Use of critics.
+We compare our method (with two critics) to variants with a single critic and to a baseline without a critic (policy-gradient).
+As shown in Appendix F.1, the two-critic architecture yields the highest accuracy, best stability, and top performance in difficult settings (e.g., cognitive), confirming the value of conservative estimation.
+Budget efficiency.
+While ACING uses a fixed 165-query budget for fair comparison with prior work, it often converges well before the budget is exhausted.
+As illustrated in Appendix F.3, many tasks reach optimal rewards within 60–80 queries, and some within 10–20, showing strong sample efficiency under constrained settings.
+White-box model.
+Using WizardLM-13B Xu et al.
+(2024) instead of Vicuna improves median test accuracy by 8 points and increases the number of best-performing tasks (Appendix F.7).
+This demonstrates that ACING can benefit from stronger decoding models, although it remains effective across architectures.
+Action dimensionality.
+We test latent action sizes $d^{\prime}\in\{5,10,20,40,100\}$.
+Results in Appendix F.5 show that $d^{\prime}=10$ and $d^{\prime}=20$ perform consistently well, while larger dimensions like $d^{\prime}=40$ yield improvements on specific tasks.
+Number of exemplars.
+Using a single exemplar performs surprisingly well and matches the 5-exemplar setup on several tasks (e.g., phonetics and summation).
+Still, five exemplars offer more consistent gains on complex tasks (Section˜F.6).
+Budget splitting.
+We explore dividing the query budget into an exploration phase and a final re-ranking phase, where top prompts are re-evaluated multiple times.
+This two-phase strategy improves median test scores by 5 points and boosts the number of best-performing tasks (Section˜F.4).
+Figure 3: Ranking distributions across cognitive tasks for all algorithms, ordered increasing by median rank.
 5 Related Work
-Many early approaches such as AutoPrompt Shin et al.
+Early approaches such as AutoPrompt Shin et al.
 (2020), FluentPrompt Shi et al.
-(2023), and other soft prompt-based methods Lester et al.
+(2023), and soft prompt tuning Lester et al.
 (2021); Li and Liang (2021); Zhong et al.
-(2021) focused on optimizing prompts for white-box LLMs and cannot be directly applied to black-box LLMs.
-Moreover, several techniques have also been proposed for optimizing prompts for grey-box LLMs: BBT Sun et al.
-(2022b), BBTv2 Sun et al.
-(2022a), and clip-tuning Chai et al.
-(2022) use evolutionary algorithms for this purpose.
-However, these methods still require access to input token embeddings and output logits, making them unsuitable for settings where only query access is allowed.
-Other approaches Li and Liang (2021); Zhong et al.
-(2021) employ continuous soft prompts by fine-tuning parameters of specific input tokens, but these often generate human-unreadable prompts and are impractical for API-accessible LLMs.
-Some other methods Deng et al.
-(2022); Zhang et al.
-(2023), created for optimizing discrete prompts instead of continuous ones, can optimize discrete tokens but require access to the LLM’s output distribution, which is typically unavailable in black-box LLMs.
-Zhou et al.
-(2023) proposed generating candidate prompts through resampling without specific optimizations, while Guo et al.
-(2024) extended this model-free approach by using evolutionary algorithms.
-However, these methods usually require a large number of queries.
-Instruction optimization in black-box LLMs recently focused on searching through the space of soft prompts, then relying on open-source white-box LLMs as a means to transform these soft prompts into discrete prompts.
+(2021) rely on access to model gradients or embeddings, limiting their applicability in black-box settings.
+Grey-box methods such as BBT and BBTv2 Sun et al.
+(2022b, a) and CLIP-tuning Chai et al.
+(2022) relax these constraints by leveraging token embeddings or logits, but are still incompatible with API-only black-box models.
+Several recent works frame instruction search as a discrete optimization problem.
+RLPrompt Deng et al.
+(2022) and Tempera Zhang et al.
+(2023) use RL to identify prompts, but assume access to token-level outputs or confidence scores.
+Alternatively, sampling- and evolution-based strategies such as APE Zhou et al.
+(2023), PromptBreeder Fernando et al.
+(2024), EvoPrompt Guo et al.
+(2024), PromptWizard Agarwal et al.
+(2024), and Auto Evol-Instruct Zeng et al.
+(2024) iteratively generate and refine candidate prompts via LLM sampling or mutation.
+While effective in some cases, these methods typically rely on large candidate pools or expensive query budgets.
+Zeroth-order optimization (e.g., AIO Qi et al.
+(2025), ZOPO Hu et al.
+(2024)) estimates gradients without backpropagation, yet incurs high computational cost and query complexity.
 InstructZero Chen et al.
-(2024) learns discrete prompts for ChatGPT by searching through the continuous space of soft prompts using Bayesian optimization, which are then transformed into discrete prompts by a white-box LLM.
-INSTINCT Lin et al.
-(2024), a successor to this approach, relies on the same approach.
-However, instead of Bayesian optimization, it uses NeuralUCB Zhou et al.
-(2020), a contextual bandit (selection) algorithm that searches within a finite set of continuous actions generated using Sobol Eriksson et al.
+(2024) and INSTINCT Lin et al.
+(2024b) apply Bayesian Optimization (BO) and NeuralUCB Zhou et al.
+(2020) within a finite action space.
+In contrast, StablePrompt Kwon et al.
+(2024) uses PPO to stabilize discrete prompt learning, but fine-tune white-box LLMs.
+Our approach differs by formulating prompt optimization as a continuous-action RL problem, enabling more efficient exploration without gradient access or large models.
+Complementary strategies include exemplar selection Wu et al.
+(2024), preference-based feedback Lin et al.
+(2024a), and best-arm identification Shi et al.
 (2024).
-Our approach combines the hybrid strategy of using white-box and black-box LLMs with an actor-critic approach that operates in a stateless, infinite continuous-action space.
-TRIPLE Shi et al.
-(2024) modeled the prompt learning problem as a best-arm-identification problem.
-However, it only considered the problem as a selection task, assuming a given pool of discrete prompts.
-In contrast, our method addresses both the generation and selection of prompts.
+However, these typically rely on fixed prompt pools or require human preference labels.
+ACING jointly addresses both the generation and selection in a unified framework, yielding high-performing prompts without pool constraints or external supervision.
 6 Conclusion
-ACING presents a promising approach for learning prompts for black-box LLMs by framing it as a stateless continuous-action RL problem.
-Our method shows significant improvements over recent baseline methods.
-Evaluations across 31 datasets reveal that ACING enhances prompt quality, achieving a median score improvement of 10 percentage points and outperforming human-crafted instructions by up to 39 percentage points.
-This sets a new benchmark for automated instruction optimization in black-box LLMs.
+We present ACING, an actor-critic RL framework for prompt optimization in black-box LLMs.
+By formulating the task as a continuous-action problem, ACING enables exploration of an infinite space through an entropy-regularized policy under strict query constraints.
+It outperforms strong baselines and surpasses human-written instructions, without any per-task tuning, demonstrating effectiveness, efficiency, and practicality.
 7 Limitations
-While our proposed results were achieved using default and commonly used values for hyperparameters in network design, learning rates, and following previous works for some common parameters, the results presented depend on these choices.
-These results could be further improved with fine-tuning, as demonstrated in the ablation studies (§ 5.4).
-Another limitation of this work, similar to previous studies, is the risk of overfitting when dealing with a limited number of validation examples.
-This can result in a prompt that minimizes the loss on the validation set but does not generalize well to the test dataset, leading to a discrepancy between training and test performance.
-Another limitation of this work, similar to previous work (Lin et al., 2024; Chen et al., 2024), is the reliance on a white-box model, where the choice of model may lead to performance variations, as demonstrated in § 5.4.2.
-Avoiding this limitation requires acting directly on the black-box LLM, which involves handling actions in a large discrete action space.
-This necessitates adapting current RL approaches for large discrete action spaces, as studied in Fourati et al.
-(2024) and the references therein.
-Appendix A Experimental Details
-A.1 Hyperparameters Remark
+While ACING achieves strong and consistent performance across diverse tasks, we do not claim universal superiority on every individual task.
+Given the stochastic and non-convex nature of prompt optimization—particularly in black-box LLMs under tight query budgets—strong prompts can occasionally arise even from random search or other optimization approaches.
+However, ACING substantially improves the chances of discovering such prompts through a principled exploration strategy grounded in actor-critic learning.
+Similar to previous prior works (Lin et al., 2024b; Chen et al., 2024; Hu et al., 2024), our reliance on a white-box model introduces variability, as model selection affects performance (Appendix F.7).
+Addressing this limitation would require operating directly on black-box LLMs, which poses challenges due to the large discrete action space.
+Adapting RL methods for such spaces, as explored in Fourati et al.
+(2024), and proposing hybrid approaches combining soft-prompt and discrete optimization, represent promising directions for future research.
+Despite the strong performance achieved with a fixed hyperparameter configuration across all tasks, we acknowledge that domain-specific tuning might yield marginal gains, particularly in edge cases.
+That said, our goal was to emphasize generality and simplicity without introducing per-task overhead—an important factor for real-world usability.
+8 Ethical Considerations
+ACING automates prompt optimization for LLMs, which can reduce manual effort but may also amplify risks associated with bias, misuse, or harmful content generation.
+We emphasize that ACING is a general optimization framework and does not guarantee ethical outputs from the underlying LLMs.
+It should be paired with appropriate content filters and safety mechanisms when deployed in sensitive domains.
+Optimized prompts could reinforce or amplify existing biases present in the base LLMs.
+Future work should incorporate fairness-aware reward functions or post-hoc bias mitigation.
+Lastly, our experiments assume legal and ethical API usage, and we caution against applying instruction optimization to restricted or proprietary models without adherence to usage policies and terms of service.
+Appendix A Table of Notations
+For clarity, we summarize all symbols used in our actor-critic framework in Table 6.
+Symbol Meaning
+$f$ Black-box LLM (only accessible via API)
+$h$ White-box LLM (frozen, used as decoder)
+$D$ Task distribution over input–output pairs $(x,y)$
+$V$ Validation set $\{(x_{j},y_{j})\}_{j=1}^{m}$
+$T$ Test set $\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{m^{\prime}}$
+$q(\cdot,\cdot)$ Scoring function measuring output quality
+$\tau$ Discrete prompt
+$\tau(z)$ Prompt generated from soft prompt $z$
+$\tau(Pa)$ Prompt from projected action vector $a$
+$\tau^{\star}$ Optimal prompt
+$z\in\mathbb{R}^{d}$ Soft prompt vector (continuous)
+$a\in[0,1]^{d^{\prime}}$ Action (latent prompt representation)
+$P\in\mathbb{R}^{d\times d^{\prime}}$ Random projection matrix
+$\pi(\cdot;\theta)$ Policy network (actor) with parameters $\theta$
+$Q_{w_{1}}(a),Q_{w_{2}}(a)$ Twin critic networks with parameters $w_{1},w_{2}$
+$Q(a)$ Value estimate: $Q(a)=\min\{Q_{w_{1}}(a),Q_{w_{2}}(a)\}$
+$J_{Q}(w)$ Critic loss: $\mathbb{E}_{a\sim D}\left[\tfrac{1}{2}(Q_{w}(a)-r(a))^{2}\right]$
+$J_{\pi}(\theta)$ Actor loss (entropy-regularized policy objective)
+$r(a)$ Reward obtained from the environment
+$\alpha$ Temperature coefficient for entropy regularization
+$H_{\text{target}}$ Target entropy for adaptive $\alpha$
+Table 6: Summary of notations used in the ACING actor-critic framework.
+Appendix B Experimental Details
+B.1 Hyperparameters
 Across the diverse tasks, in the main paper, the same hyperparameters were used, which shows that the algorithm generalizes well across the 30 tasks without specifically tuning hyperparameters in each task.
-A summary of the key parameters cab be found in the following Table.
-An implementation of ACING is available at https://github.com/salmakh1/ACING.
+A summary of the key parameters can be found in the following Table.
 Hyperparameter Choice
-White-box $h$ Vicuna-13B
-Actor-network $(1-1024-256-10)$
-Critic-network $(10-128-128-1)$
+White-box $h$ Vicuna-13B and WizardLM
+Actor-network (1102425610)(1-1024-256-10)
+Critic-network (101281281)(10-128-128-1)
 Budget $T$ 165
-Intrinsic (action) dimension $d^{\prime}$ 5
+Intrinsic (action) dimension $d^{\prime}$ 10
 Number of soft tokens $N_{z}$ 5
@@
 Table 7: Key hyperparameters and their values.
-A.2 Actor-Critic Details
-Across all the tasks, we used three fully-connected layers for both the actor $(1-1024-256-10)$ and the critics $(10-128-128-1)$ networks, with learning rates fixed at $3\cdot 10^{-4}$ for each.
-We used two critic networks and consider their minimum as the actual critic.
+Furthermore, like previous works, we use their default (tuned) hyperparameters for the results in the main paper, including the intrinsic dimension $d^{\prime}=10$ and the number of soft tokens $N_{z}=5$.
+For fairness, we refrain from fine-tuning these parameters for our method and use the same values as in prior works.
+This ensures that our ACING algorithm searches in the same space, $0,1]^{10}$, and uses the same total number of queries to the black-box LLM as APE, InstructZero, and INSTINCT for a fair comparison.
+For each algorithm, after identifying the best instruction using the validation set $\mathcal{V}$, we evaluate the discovered instruction on a separate test set $\mathcal{T}$ and report the test score.
+B.2 Actor-Critic Details
+Across all the tasks, we used three fully-connected layers for both the actor (1102425610)(1-1024-256-10) and the critics (101281281)(10-128-128-1) networks, with learning rates fixed at $3\cdot 10^{-4}$ for each.
 We learn the entropy parameter $\alpha$ using a learning rate of $9\cdot 10^{-4}$.
-Appendix B ACING Rewards over the (Calls) Steps
+We adopt two independently trained critics and take the minimum of their outputs to mitigate overestimation bias.
+This conservative approach helps regularize training and improves robustness, particularly in tasks with noisy or sparse rewards.
+We further validate this choice in our ablation studies with one critic, two critics, and without critics.
+B.3 Metrics
+Table 8 outlines the evaluation metrics used across various task types.
+Depending on the nature of the task, we adopt different scoring schemes to ensure fair and meaningful evaluation.
+For most classification and generation tasks, we employ exact match (EM) scoring, which requires the prediction to match the ground truth exactly, making it a stringent yet interpretable metric.
+B.4 Licenses and Terms of Use for Artifacts
+Black-box APIs.
+We use the OpenAI API to access ChatGPT OpenAI (2023a), GPT-4 OpenAI (2023c), and GPT-4o Hurst et al.
+(2024) for black-box evaluation.
+These models are proprietary and accessed via paid API under OpenAI’s terms of service.
+We do not redistribute or modify these models.
+White-box Models.
+Our framework uses Vicuna-13B Chiang et al.
+(2023) (and WizardLM Xu et al.
+(2024) for ablations) as a frozen decoder to generate discrete prompts.
+Both models are publicly available for research use under a non-commercial license.
+We follow all usage restrictions and do not modify or redistribute these models.
+Datasets.
+We use publicly available datasets including:
+  • Instruction induction datasets from Honovich et al. (2023).
+  • SAMSum (Gliwa et al., 2019) for summarization.
+  • GSM8K (Cobbe et al., 2021) and AQUA-RAT (Ling et al., 2017) for reasoning.
+All datasets are used under their original licenses and are cited appropriately.
+Our Code and Models.
+We will publicly release our implementation, including the actor-critic training framework and prompt optimization pipeline, under an open-source license (MIT license) upon publication.
+B.5 Data Usage, Privacy, and Safety Considerations
+All datasets for the studied tasks are publicly released and widely used in NLP research. We used them under their respective licenses for academic use, and we cite the original sources in all cases.
+Privacy and Anonymization.
+None of the datasets we used contain personally identifiable information (PII) to the best of our knowledge.
+The datasets are either synthetic, anonymized, or derived from public sources with appropriate preprocessing.
+We did not augment or modify these datasets in ways that would introduce identifying information.
+Human Evaluation and Safety.
+For our human evaluation of ACING-generated prompts, we used only automatically generated instructions and task templates.
+Prior to annotation, we reviewed the generated prompts and outputs to ensure that they did not contain any offensive or inappropriate content.
+The tasks involved abstract or anonymized content (e.g., generic input strings or public task descriptions), and no human names, images, or personal data were included.
+Ethical Use.
+We ensured that all experiments involving human subjects were conducted ethically and responsibly.
+Participation in the annotation study was fully voluntary.
+Annotators were informed about the nature and purpose of the task and could withdraw at any time.
+The study was limited to rating task-level clarity and alignment, involved no sensitive content, and did not collect any personal or identifiable information.
+All responses were anonymized, and only aggregate statistics were analyzed.
+As the study posed no foreseeable risk to participants and did not involve identifiable data, it did not require formal ethics review under our institution’s guidelines.
+B.6 Artifact Documentation
+Task Coverage.
+We evaluate ACING on a diverse set of NLP tasks, categorized as follows:
+  • Instruction Induction: A wide range of classification and transformation tasks drawn from Honovich et al. (2023), covering syntax, semantics, phonetics, word manipulation, and reasoning.
+  • Reasoning: Zero-shot chain-of-thought reasoning tasks using GSM8K (Cobbe et al., 2021) and AQUA-RAT (Ling et al., 2017), which involve multi-step symbolic and arithmetic problem solving.
+  • Summarization: Dialogue-based abstractive summarization using the SAMSum dataset (Gliwa et al., 2019).
+  • Translation: Lexical-level English-to-German, English-to-Spanish, and English-to-French tasks to assess multilingual instruction learning.
+Language Coverage.
+Our experiments involve four languages: English (the primary language for most tasks), and German, Spanish, and French in translation tasks.
+Linguistic Phenomena.
+The tasks cover a broad range of linguistic phenomena, including:
+  • Morphosyntax: e.g., negation, word reordering.
+  • Lexical Semantics: e.g., antonyms, synonyms, word similarity.
+  • Discourse and Pragmatics: e.g., summarization, cause-effect reasoning.
+  • Symbolic and Numerical Reasoning: e.g., math problem solving, program induction.
+Demographics and Bias.
+All datasets used are publicly available academic benchmarks.
+None were collected or annotated by the authors.
+To the best of our knowledge, these datasets do not contain personally identifiable information (PII) or explicit demographic attributes.
+No demographic inference, fairness evaluation, or bias analysis was performed, as our work focuses on instruction optimization and not on social or identity-based NLP tasks.
+Human Annotation.
+We conducted a human evaluation study to assess the clarity and semantic alignment of instructions generated by ACING.
+Annotators were presented with anonymized task prompts and model-generated instructions, and asked to rate clarity and alignment on a 5-point Likert scale.
+The annotation involved 26 in-house participants, all of whom were informed of the task goals and free to opt out at any time.
+No demographic data was collected.
+All instructions shown were manually reviewed to ensure that they contained no PII or offensive content.
+Collected responses were fully anonymized, and no free-form sensitive text was stored.
+Results are summarized in Section 4.3 and Appendix˜H.
+The full set of instructions used in the study is from Table˜2.
+Artifact Availability.
+We will release our full codebase under an open-source license (MIT license) upon publication.
+This includes the reinforcement learning framework, prompt decoding pipeline, and evaluation scripts.
+All third-party datasets and models used are cited and publicly available under their respective research licenses.
+B.7 Implementation and Compute Details
+Model Sizes.
+Our actor and critic networks are lightweight MLPs with three layers each.
+The white-box decoder is Vicuna-13B, and the black-box models queried include GPT-3.5 (ChatGPT), GPT-4, and GPT-4o, whose sizes are not publicly disclosed.
+Compute Infrastructure.
+We use an internal SLURM cluster for running our experiments.
+The experiments were done on an ASUS ESC N4A-E11 server.
+The node has 4 A100 GPUs, an AMD EPYC 7003 series 64 core @ 3.5GHz CPU and 512GB of RAM.
+We used one A100, with 2 cores, and required at most 50GB of memory for the experiments.
+All black-box LLM queries were performed via API.
+Compute Budget.
+Each ACING run used a fixed query budget of 165 black-box API calls.
+Across 33 tasks and 3 random seeds, this corresponds to a total of approximately 16,000 black-box LLM queries.
+Training time per task was approximately 10–20 minutes on a single GPU.
+Task Score Type
+Common_concept, Informal_to_formal F1 score = $(2\times\text{precision}\times\text{recall})/(\text{precision}+\text{recall})$
+Orthography_starts_with, Taxonomy_animal Set match: prediction set must match ground truth set
+Synonyms In-list match: prediction is correct if it falls within the list of ground truth words
+All other datasets (e.g., Antonyms, Translation, Cause_and_effect, Diff) EM score: 1 if exact match, 0 otherwise (letter-by-letter for words)
+Table 8: Scoring metrics for different tasks.
+Appendix C ACING vs. Other Optimization Methods
+We compare our method against recent baselines on the 30 instruction-induction datasets.
+The results in Table 9 show the average test accuracy (along with the standard deviation) over three independent runs, using three different seeds.
+For each seed, we selected the best instruction achieved by each method and evaluated it on the testing dataset.
+The table demonstrates that our method, ACING, outperforms others by achieving the highest accuracy in 14 out of 30 tasks, compared to INSTINCT, InstructZERO, EvoPrompt, and APE which succeeded in 8 tasks or less each.
+Additionally, ACING achieves the highest median accuracy across tasks, with a value of 0.71, which is  22 percentage points higher than APE.
+Table 21 shows the best prompt achieved for each task and corresponding test scores.
+Category Task APE EvoPrompt InstructZero INSTINCT ACING
+Spelling Letters_list 0.59 (0.02) 0.97 (0.03) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00)
+First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
+Second_word_letter 0.00 (0.00) 0.63 (0.17) 0.35 (0.09) 0.39 (0.28) 0.70 (0.15)
+Morpho-Syntax Singular_to_plural 1.00 (0.00) 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.95 (0.03)
+Active_to_passive 1.00 (0.00) 0.99 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00)
+Negation 0.79 (0.00) 0.84 (0.02) 0.65 (0.10) 0.58 (0.22) 0.71 (0.06)
+Lexical Semantics Antonyms 0.79 (0.02) 0.70 (0.01) 0.76 (0.00) 0.84 (0.01) 0.74 (0.01)
+Synonyms 0.14 (0.01) 0.19 (0.07) 0.22 (0.11) 0.19 (0.08) 0.13 (0.02)
+Word_unscrambling 0.54 (0.00) 0.44 (0.06) 0.59 (0.06) 0.54 (0.02) 0.50 (0.07)
+Phonetics Rhymes 0.59 (0.01) 0.52 (0.05) 0.99 (0.01) 0.36 (0.04) 0.57 (0.31)
+Numerical Sum 0.87 (0.01) 1.00 (0.00) 1.00 (0.00) 0.70 (0.21) 1.00 (0.00)
+Diff 0.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.93 (0.09) 1.00 (0.00)
+Knowledge Larger_animal 0.72 (0.02) 0.58 (0.06) 0.63 (0.07) 0.81 (0.09) 0.84 (0.07)
+Periodic_elements 0.99 (0.01) 0.92 (0.00) 0.96 (0.03) 1.00 (0.00) 0.98 (0.00)
+Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.48 (0.10) 0.52 (0.09) 0.55 (0.11) 0.69 (0.15)
+Common_concept 0.03 (0.02) 0.17 (0.00) 0.14 (0.04) 0.09 (0.04) 0.19 (0.05)
+Object_counting 0.30 (0.02) 0.50 (0.06) 0.38 (0.06) 0.40 (0.12) 0.41 (0.03)
+Odd_one_out 0.32 (0.02) 0.64 (0.04) 0.57 (0.02) 0.25 (0.18) 0.64 (0.00)
+Orthography_starts_with 0.23 (0.01) 0.47 (0.02) 0.41 (0.09) 0.54 (0.06) 0.60 (0.12)
+Taxonomy_animal 0.02 (0.02) 0.38 (0.15) 0.67 (0.14) 0.85 (0.06) 0.71 (0.02)
+Auto_categorization 0.31 (0.01) 0.20 (0.03) 0.29 (0.02) 0.07 (0.07) 0.29 (0.04)
+Word_sorting 0.58 (0.01) 0.01 (0.00) 0.64 (0.05) 0.23 (0.20) 0.70 (0.03)
+CLUE Sentence_similarity 0.00 (0.00) 0.05 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.07)
+Sentiment 0.90 (0.00) 0.63 (0.17) 0.88 (0.03) 0.88 (0.02) 0.89 (0.01)
+Translation Num_to_verbal 0.13 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.99 (0.01)
+Translation_en-de 0.83 (0.01) 0.80 (0.02) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01)
+Translation_en-es 0.86 (0.01) 0.76 (0.00) 0.67 (0.24) 0.89 (0.00) 0.87 (0.02)
+Translation_en-fr 0.88 (0.01) 0.86 (0.00) 0.77 (0.06) 0.85 (0.02) 0.83 (0.01)
+Style Informal_to_formal 0.57 (0.01) 0.50 (0.02) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05)
+Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.25 (0.00)
+median score 0.49 0.63 0.66 0.64 0.71
+# best-performing tasks 8 8 8 7 14
+Table 9: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE (Zhou et al., 2023), EvoPrompt (Guo et al., 2024), InstructZero (Chen et al., 2024), and INSTINCT Lin et al.
+(2024b).
+The bottom rows report the median score and total number of best-performing tasks for each method.
+Appendix D ACING vs InstructZero on GPT-4
+We extend our experiments to include GPT-4 as the black-box LLM and Vicuna as the white-box model.
+Given the high cost associated with querying GPT-4, we restrict our comparison to InstructZero, the strongest baseline after ACING based on prior results with GPT-3.5 (Table 9).
+To ensure a fair and cost-efficient comparison, both methods are allocated an equal budget of 100 API calls.
+Additionally, we focus this analysis on the most challenging subset—cognitive tasks—to concentrate the evaluation on settings where prompt optimization is most demanding.
+Table 10 presents the results of this experiment.
+Consistent with our earlier findings (Table 9), ACING continues to excel in cognitive tasks, solving 6 out of 8 tasks compared to only 3 solved by InstructZero, and achieving a 13-point higher median score.
+These results further highlight ACING’s robustness and effectiveness under tighter budgets and more difficult evaluation scenarios.
+Category Task InstructZero ACING
+Cognitive Tasks on GPT-4 Cause_and_effect 0.43 (0.10) 0.53 (0.10)
+Common_concept 0.27 (0.00) 0.04 (0.01)
+Object_counting 0.54 (0.03) 0.62 (0.03)
+Odd_one_out 0.77 (0.01) 0.77 (0.01)
+Orthography_starts_with 0.40 (0.12) 0.60 (0.03)
+Taxonomy_animal 0.97 (0.02) 0.98 (0.00)
+Auto_categorization 0.35 (0.01) 0.31 (0.02)
+Word_sorting 0.72 (0.03) 0.74 (0.01)
+median score 0.48 0.61
+# best-performing tasks 3 6
+Table 10: Average test performance (with standard deviations) across 3 random seeds comparing ACING and InstructZero Chen et al.
+(2024) on 8 cognitively demanding tasks using GPT-4 (black-box) and Vicuna (white-box), under a budget of 100 calls.
+The bottom rows report the median score and the total number of tasks where each method achieves the highest performance.
+Appendix E Extended Results on GPT-4o
+We extend a subset of our evaluation to include results on GPT-4o Hurst et al.
+(2024), as shown in Table 11.
+These results are based on 3 repetitions, and we report the average accuracy along with standard deviation in parentheses.
+We observe that ACING continues to perform competitively, often outperforming human-written instructions on tasks such as Antonyms, Rhyme, and Sentence Similarity.
+Task with GPT-4o Human ACING
+Antonyms 0.73 (0.01) 0.80 (0.01)
+Rhyme 0.61 (0.01) 0.80 (0.15)
+Sentence Similarity 0.00 (0.02) 0.15 (0.00)
+Informal to Formal 0.58 (0.00) 0.48 (0.02)
+Synonyms 0.14 (0.00) 0.20 (0.05)
+Cause and Effect 0.92 (0.03) 0.85 (0.13)
+Table 11: Comparison of human-written (from Table˜19 vs.
+ACING-generated instructions on a subset of tasks evaluated using GPT-4o Hurst et al.
+(2024).
+Results are averaged over 3 repetitions; standard deviations are shown in parentheses.
+Appendix F Ablation Studies
+We perform ablation studies to understand the role of key design choices in ACING’s performance.
+These include critic usage, action dimensionality, exemplar count, budget allocation, and white-box model selection.
+F.1 On the use of Critics
+To empirically assess the contribution of the critic(s), we evaluate three variants of our method:
+  • ACING (two critics): Full method using dual critics.
+  • ACING (one critic): Ablation using a single critic.
+  • Policy Gradient (no critic): Baseline using pure policy gradients without any critic.
+We report results averaged across 30 diverse tasks (3 trials per task). Due to space constraints, we summarize performance using per-category averages and global statistics.
+Summary: In Table 12, ACING achieves a slightly higher median score (0.71 vs.
+0.70) and clearly dominates in terms of robustness, outperforming policy gradient on 21 of 33 tasks.
+This suggests that the critic plays a crucial role in stabilizing learning and guiding exploration, especially in noisy or sparse reward settings (e.g., Rhymes, Auto_categorization, Word_sorting).
+Table 13 further underscores the benefit of using two critics: while both ACING variants reach the same median score (0.71), the two-critic version leads on more tasks (22 vs.
+16) and shows stronger performance on several cognitively demanding or unstable tasks (e.g., Cause_and_effect, Sentence_similarity, Taxonomy_animal).
+This highlights the importance of ensemble value estimation in improving reliability across diverse tasks.
+Category Task ACING Policy Gradient (no critic)
+Spelling Letters_list 1.00 (0.00) 0.93 (0.05)
+First_word_letter 1.00 (0.00) 1.00 (0.00)
+Second_word_letter 0.70 (0.15) 0.55 (0.30)
+Morpho-Syntax Singular_to_plural 0.95 (0.03) 1.00 (0.00)
+Active_to_passive 1.00 (0.00) 1.00 (0.00)
+Negation 0.71 (0.06) 0.71 (0.00)
+Lexical Semantics Antonyms 0.74 (0.01) 0.69 (0.15)
+Synonyms 0.13 (0.02) 0.19 (0.10)
+Word_unscrambling 0.50 (0.07) 0.54 (0.02)
+Phonetics Rhymes 0.57 (0.31) 0.36 (0.14)
+Numerical Sum 1.00 (0.00) 0.98 (0.03)
+Diff 1.00 (0.00) 0.93 (0.05)
+Knowledge Larger_animal 0.84 (0.07) 0.68 (0.18)
+Periodic_elements 0.98 (0.00) 0.93 (0.04)
+Cognitive Tasks Cause_and_effect 0.69 (0.15) 0.71 (0.19)
+Common_concept 0.19 (0.05) 0.12 (0.02)
+Object_counting 0.41 (0.03) 0.44 (0.08)
+Odd_one_out 0.64 (0.00) 0.53 (0.08)
+Orthography_starts_with 0.60 (0.12) 0.52 (0.14)
+Taxonomy_animal 0.71 (0.02) 0.77 (0.10)
+Auto_categorization 0.29 (0.04) 0.17 (0.12)
+Word_sorting 0.70 (0.03) 0.23 (0.30)
+CLUE Sentence_similarity 0.13 (0.07) 0.00 (0.00)
+Sentiment 0.89 (0.01) 0.89 (0.00)
+Translation Num_to_verbal 0.99 (0.01) 1.00 (0.00)
+Translation_en-de 0.82 (0.01) 0.80 (0.00)
+Translation_en-es 0.87 (0.02) 0.87 (0.01)
+Translation_en-fr 0.83 (0.01) 0.87 (0.00)
+Style Informal_to_formal 0.44 (0.05) 0.50 (0.03)
+Coding Auto_debugging 0.25 (0.00) 0.25 (0.00)
+median score 0.71 0.70
+# best-performing tasks 21 14
+Table 12: Performance comparison between ACING and Reinforce different task categories.
+Category Task ACING (two critics) ACING (one critic)
+Spelling Letters_list 1.00 (0.00) 1.00 (0.00)
+First_word_letter 1.00 (0.00) 1.00 (0.00)
+Second_word_letter 0.70 (0.15) 0.58 (0.31)
+Morpho-Syntax Singular_to_plural 0.95 (0.03) 0.99 (0.00)
+Active_to_passive 1.00 (0.00) 1.00 (0.00)
+Negation 0.71 (0.06) 0.78 (0.01)
+Lexical Semantics Antonyms 0.74 (0.01) 0.72 (0.05)
+Synonyms 0.13 (0.02) 0.13 (0.00)
+Word_unscrambling 0.50 (0.07) 0.52 (0.04)
+Phonetics Rhymes 0.57 (0.31) 0.63 (0.23)
+Numerical Sum 1.00 (0.00) 0.98 (0.01)
+Diff 1.00 (0.00) 0.99 (0.02)
+Knowledge Larger_animal 0.84 (0.07) 0.80 (0.12)
+Periodic_elements 0.98 (0.00) 0.97 (0.02)
+Cognitive Tasks Cause_and_effect 0.69 (0.15) 0.56 (0.00)
+Common_concept 0.19 (0.05) 0.16 (0.09)
+Object_counting 0.41 (0.03) 0.44 (0.09)
+Odd_one_out 0.64 (0.00) 0.59 (0.05)
+Orthography_starts_with 0.60 (0.12) 0.63 (0.12)
+Taxonomy_animal 0.71 (0.02) 0.59 (0.40)
+Auto_categorization 0.29 (0.04) 0.29 (0.06)
+Word_sorting 0.70 (0.03) 0.70 (0.01)
+CLUE Sentence_similarity 0.13 (0.07) 0.05 (0.05)
+Sentiment 0.89 (0.01) 0.88 (0.02)
+Translation Num_to_verbal 0.99 (0.01) 1.00 (0.00)
+Translation_en-de 0.82 (0.01) 0.81 (0.01)
+Translation_en-es 0.87 (0.02) 0.87 (0.02)
+Translation_en-fr 0.83 (0.01) 0.82 (0.02)
+Style Informal_to_formal 0.44 (0.05) 0.49 (0.05)
+Coding Auto_debugging 0.25 (0.00) 0.25 (0.00)
+median score 0.71 0.71
+# best-performing tasks 22 16
+Table 13: Performance comparison between ACING, and ACING_one_critic across different task categories.
+F.2 On Direct Parameterization of the Actor
+To assess the impact of using a feed-forward network (FFN) versus direct parameterization for the actor, we compare the following two variants:
+  • ACING: Our standard approach, where the actor is parameterized using a feed-forward neural network that receives a constant one-dimensional input (set to 1).
+  • ACING (direct parameterization): A variant using the same critics, but with the actor represented directly by learnable mean and variance vectors, completely removing the FFN.
+We report results averaged across 30 diverse tasks (3 trials per task), summarizing performance by task category and global statistics.
+Summary: As shown in Table 14, the FFN-based ACING achieves a substantially higher median score (0.69 vs.
+0.44) and outperforms the direct parameterization variant on many more tasks (19 vs.
+8).
+While direct parameterization occasionally matches or surpasses FFN performance on simpler tasks (e.g., Negation, Antonyms, Num_to_verbal, Informal_to_formal), it generally struggles on tasks requiring more compositional or fine-grained control (e.g., Word_sorting, Odd_one_out, Auto_categorization).
+Category Task ACING ACING (w/o FFN)
+Spelling Letters_list 1.00 (0.00) 0.99 (0.00)
+First_word_letter 1.00 (0.00) 0.97 (0.03)
+Second_word_letter 0.70 (0.15) 0.21 (0.51)
+Negation 0.71 (0.06) 0.81 (0.03)
+Lexical Semantics Antonyms 0.74 (0.01) 0.75 (0.06)
+Synonyms 0.13 (0.02) 0.25 (0.09)
+Word_unscrambling 0.50 (0.07) 0.41 (0.06)
+Phonetics Rhymes 0.57 (0.31) 0.38 (0.14)
+Numerical Sum 1.00 (0.00) 0.79 (0.15)
+Diff 1.00 (0.00) 0.86 (0.12)
+Knowledge Larger_animal 0.84 (0.07) 0.76 (0.05)
+Cognitive Tasks Cause_and_effect 0.69 (0.15) 0.53 (0.07)
+Common_concept 0.19 (0.05) 0.11 (0.01)
+Object_counting 0.41 (0.03) 0.33 (0.08)
+Odd_one_out 0.64 (0.00) 0.37 (0.20)
+Orthography_starts_with 0.60 (0.12) 0.39 (0.17)
+Taxonomy_animal 0.71 (0.02) 0.67 (0.06)
+Auto_categorization 0.29 (0.04) 0.11 (0.15)
+Word_sorting 0.70 (0.03) 0.19 (0.26)
+CLUE Sentence_similarity 0.13 (0.07) 0.00 (0.00)
+Translation Num_to_verbal 0.99 (0.01) 1.00 (0.00)
+Translation_en-es 0.87 (0.02) 0.87 (0.01)
+Style Informal_to_formal 0.44 (0.05) 0.48 (0.03)
+Coding Auto_debugging 0.25 (0.00) 0.25 (0.00)
+median score 0.69 0.44
+# best-performing tasks 19 7
+Table 14: Performance comparison between ACING and a variant using the same critics but with direct parameterization of the actor (learnable mean and variance vectors) instead of a feed-forward network.
+F.3 ACING Rewards over the (Calls) Steps
 In the main paper, we report the final test score after a fixed budget of 165 black-box LLM calls.
@@
 The choice of 165 calls was mainly based on previous work Lin et al.
-(2024); Chen et al.
+(2024b); Chen et al.
 (2024), avoiding any potential advantage that could come from optimizing this number.
@@
 Figure 4: Reward plots for running ACING on various selected tasks, showing the highest achieved reward on the y-axis until each API call (step), with the x-axis representing the number of API calls.
-Appendix C ACING with Budget Splitting
+F.4 ACING with Budget Splitting
 Due to the stochastic nature of the black-box LLM, the same instruction may yield different rewards when evaluated by the LLM.
@@
 Finally, the prompt with the highest average reward across repetitions is used at test time.
-In Table 5, we demonstrate that ACING, with an exploration budget of $T=150$ and the remaining 15 calls allocated to uniform exploration of the top $p=5$ prompts (evaluated $k=3$ times each), achieves improved median scores across tasks and higher test accuracy on 15 tasks compared to previous work, unlike without the splitting technique which achieved higher test accuracy on 14 tasks compared to the previous work.
-Category Task APE InstructZero INSTINCT ACING150+15 ACING165
+In Table 15, we demonstrate that ACING, with an exploration budget of $T=150$ and the remaining 15 calls allocated to uniform exploration of the top $p=5$ prompts (evaluated $k=3$ times each), achieves improved median scores across tasks and higher test accuracy on 13 tasks compared to previous work.
+Furthermore, it acheives a higher median score compared to ACING without splitting.
+Category Task APE EvoPrompt InstructZero INSTINCT ACING150+15 ACING165
 (budget splitting) (main paper)
-Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00)
-First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
-Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.40 (0.17) 0.70 (0.15)
-Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.99 (0.01) 0.95 (0.03)
-Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
-Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.82 (0.00) 0.71 (0.06)
-Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.76 (0.06) 0.74 (0.01)
-Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.12 (0.02) 0.13 (0.02)
-Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.59 (0.05) 0.50 (0.07)
-Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.60 (0.38) 0.57 (0.31)
-Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.81 (0.30) 0.69 (0.31)
-Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 0.97 (0.04) 1.00 (0.00)
-Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.86 (0.06) 0.84 (0.07)
-Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.03) 0.98 (0.00)
-Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.76 (0.18) 0.69 (0.15)
-Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.10 (0.01) 0.19 (0.05)
-Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.48 (0.11) 0.41 (0.03)
-Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.59 (0.05) 0.64 (0.00)
-Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.54 (0.15) 0.60 (0.12)
-Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.53 (0.34) 0.71 (0.02)
-Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.27 (0.06) 0.29 (0.04)
-Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.72 (0.02) 0.70 (0.03)
-CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.08) 0.13 (0.07)
-Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.88 (0.03) 0.89 (0.01)
-Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
-Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01) 0.82 (0.01)
-Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.86 (0.02) 0.87 (0.02)
-Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.85 (0.02) 0.83 (0.01)
-Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05) 0.44 (0.05)
-Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.29 (0.07) 0.25 (0.00)
-median score 0.49 0.66 0.64 0.76 0.71
-# best-performing tasks 7 8 8 15 14
-Table 5: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT.
+Spelling Letters_list 0.59 (0.02) 0.97 (0.03) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00)
+First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
+Second_word_letter 0.00 (0.00) 0.63 (0.17) 0.35 (0.09) 0.39 (0.28) 0.40 (0.17) 0.70 (0.15)
+Morpho-Syntax Singular_to_plural 1.00 (0.00) 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.99 (0.01) 0.95 (0.03)
+Active_to_passive 1.00 (0.00) 0.99 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
+Negation 0.79 (0.00) 0.84 (0.02) 0.65 (0.10) 0.58 (0.22) 0.82 (0.00) 0.71 (0.06)
+Lexical Semantics Antonyms 0.79 (0.02) 0.70 (0.01) 0.76 (0.00) 0.84 (0.01) 0.76 (0.06) 0.74 (0.01)
+Synonyms 0.14 (0.01) 0.19 (0.07) 0.22 (0.11) 0.19 (0.08) 0.12 (0.02) 0.13 (0.02)
+Word_unscrambling 0.54 (0.00) 0.44 (0.06) 0.59 (0.06) 0.54 (0.02) 0.59 (0.05) 0.50 (0.07)
+Phonetics Rhymes 0.59 (0.01) 0.52 (0.05) 0.99 (0.01) 0.36 (0.04) 0.60 (0.38) 0.57 (0.31)
+Numerical Sum 0.87 (0.01) 1.00 (0.00) 1.00 (0.00) 0.70 (0.21) 0.98 (0.01) 1.00 (0.00)
+Diff 0.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.93 (0.09) 0.97 (0.04) 1.00 (0.00)
+Knowledge Larger_animal 0.72 (0.02) 0.58 (0.06) 0.63 (0.07) 0.81 (0.09) 0.86 (0.06) 0.84 (0.07)
+Periodic_elements 0.99 (0.01) 0.92 (0.00) 0.96 (0.03) 1.00 (0.00) 0.98 (0.03) 0.98 (0.00)
+Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.48 (0.10) 0.52 (0.09) 0.55 (0.11) 0.76 (0.18) 0.69 (0.15)
+Common_concept 0.03 (0.02) 0.17 (0.00) 0.14 (0.04) 0.09 (0.04) 0.10 (0.01) 0.19 (0.05)
+Object_counting 0.30 (0.02) 0.50 (0.06) 0.38 (0.06) 0.40 (0.12) 0.48 (0.11) 0.41 (0.03)
+Odd_one_out 0.32 (0.02) 0.64 (0.04) 0.57 (0.02) 0.25 (0.18) 0.59 (0.05) 0.64 (0.00)
+Orthography_starts_with 0.23 (0.01) 0.47 (0.02) 0.41 (0.09) 0.54 (0.06) 0.54 (0.15) 0.60 (0.12)
+Taxonomy_animal 0.02 (0.02) 0.38 (0.06) 0.67 (0.14) 0.85 (0.06) 0.53 (0.34) 0.71 (0.02)
+Auto_categorization 0.31 (0.01) 0.20 (0.03) 0.29 (0.02) 0.07 (0.07) 0.27 (0.06) 0.29 (0.04)
+Word_sorting 0.58 (0.01) 0.01 (0.00) 0.64 (0.05) 0.23 (0.20) 0.72 (0.02) 0.70 (0.03)
+CLUE Sentence_similarity 0.00 (0.00) 0.05 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.08) 0.13 (0.07)
+Sentiment 0.90 (0.00) 0.63 (0.17) 0.88 (0.03) 0.88 (0.02) 0.88 (0.03) 0.89 (0.01)
+Translation Num_to_verbal 0.13 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
+Translation_en-de 0.83 (0.01) 0.80 (0.02) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01) 0.82 (0.01)
+Translation_en-es 0.86 (0.01) 0.76 (0.00) 0.67 (0.24) 0.89 (0.00) 0.86 (0.02) 0.87 (0.02)
+Translation_en-fr 0.88 (0.01) 0.86 (0.00) 0.77 (0.06) 0.85 (0.02) 0.85 (0.02) 0.83 (0.01)
+Style Informal_to_formal 0.57 (0.01) 0.50 (0.02) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05) 0.44 (0.05)
+Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.29 (0.07) 0.25 (0.00)
+median score 0.49 0.63 0.66 0.64 0.76 0.71
+# best-performing tasks 7 7 7 8 13 13
+Table 15: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, EvoPrompt, and INSTINCT.
 The bottom rows report the median score and total number of best-performing tasks for each method.
-Appendix D ACING with Different Intrinsic (Action) Dimensions
+F.5 ACING with Different Intrinsic (Action) Dimensions
 In the main paper, we present results using actions with a dimension of $d^{\prime}=10$, following the setup of prior work.
@@
 Phonetics Rhymes 0.95 0.36 0.94 1.00 1.00
-Numerical Sum 0.99 0.79 1.00 0.99 1.00
+Numerical Sum 0.99 1.00 1.00 0.99 1.00
 Diff 0.89 1.00 1.00 1.00 1.00
@@
 Coding Auto_debugging 0.25 0.25 0.25 0.25 0.25
-# best-performing tasks 8 9 10 15 14
+# best-performing tasks 8 10 10 15 14
 Table 16: Average ACING test performance for a fixed random seed (0) with different soft prompt dimensions $d^{{}^{\prime}}$. The bottom row report the total number of best-performing tasks.
-Appendix E ACING with Different Number of Exemplars
+F.6 ACING with Different Number of Exemplars
 In this section, we test ACING with a single exemplar, in contrast to the main results in the paper, which use five exemplars for ACING and all other benchmarks.
@@
 Phonetics Rhymes 0.89 (0.08) 0.57 (0.31)
-Numerical Sum 0.99 (0.01) 0.69 (0.31)
+Numerical Sum 0.99 (0.01) 1.00 (0.00)
 Diff 0.99 (0.01) 1.00 (0.00)
@@
 median score 0.67 0.71
-# best-performing tasks 12 23
+# best-performing tasks 11 24
 Table 17: Average ACING test performance (and standard deviations) across 3 random seeds comparing 1 exemplar versus 5 exemplars.
 The bottom rows report the median score and total number of best-performing tasks.
-Appendix F ACING with Different White-box models
+F.7 ACING with Different White-box models
 […]
@@
 Phonetics Rhymes 0.57 (0.31) 0.97 (0.04)
-Numerical Sum 0.69 (0.31) 1.00 (0.00)
+Numerical Sum 1.00 (0.00) 1.00 (0.00)
 Diff 1.00 (0.00) 1.00 (0.00)
@@
 median score 0.71 0.79
-# best-performing tasks 14 21
+# best-performing tasks 15 21
 Table 18: Average ACING test performance (and standard deviations) across 3 random seeds using Vicuna and WizardLM as white-box models.
@@
 Appendix G Demonstrations with Human Instructions
+To contextualize the performance of ACING, we compare its best-learned instructions against human-written instructions curated by Honovich et al.
+(2023).
+Table 19 presents a representative subset of tasks, categorized by linguistic and semantic attributes, along with input–output demonstrations, human instructions, and corresponding performance scores.
+While human instructions often perform strongly, ACING matches or exceeds them in the majority of cases, particularly on tasks like Antonyms, Rhymes, and Sentence Similarity, where learned instructions yield notable improvements.
+The comparison underscores ACING’s capacity not only to automate instruction crafting but also to outperform carefully designed human-written prompts across diverse task types.
+Summary statistics at the bottom of the table show that ACING achieves a higher average and median score, and wins on a greater number of tasks overall.
 Category Task Demonstration Human Instruction Honovich et al. (2023) Human Score Our Score
-Spelling First_word_letter cat $\rightarrow$ c Extract the first letter of the input word. 1.00 1.00
-Second_word_letter cat $\rightarrow$ a Extract the second letter of the input word. 0.96 0.59
-Letters_list cat $\rightarrow$ c a t Break the input word into letters, separated by spaces. 1.00 1.00
-Morpho- syntax Singular_to_plural cat $\rightarrow$ cats Convert the input word to its plural form. 1.00 1.00
-Active_to_passive The artist introduced the scientist. $\rightarrow$ The scientist was introduced by the artist. Write the input sentence in passive form. 1.00 1.00
-Syntax Negation Time is finite $\rightarrow$ Time is not finite. Negate the input sentence. 0.81 0.82
-Lexical Semantics Antonyms won $\rightarrow$ lost Write a word that means the opposite of the input word. 0.70 0.82
-Synonyms alleged $\rightarrow$ supposed Write a word with a similar meaning to the input word. 0.14 0.13
-Phonetics Rhymes sing $\rightarrow$ ring Write a word that rhymes with the input word. 0.61 1.00
-Knowledge Larger_animal koala, snail $\rightarrow$ koala Write the larger of the two given animals. 0.94 0.94
+Spelling First_word_letter cat $\rightarrow$ c Extract the first letter of the input word. 1.00 (0.00) 1.00 (0.00)
+Second_word_letter cat $\rightarrow$ a Extract the second letter of the input word. 0.96 (0.00) 0.92 (0.00)
+Letters_list cat $\rightarrow$ c a t Break the input word into letters, separated by spaces. 1.00 (0.00) 1.00 (0.00)
+Morpho- syntax Singular_to_plural cat $\rightarrow$ cats Convert the input word to its plural form. 1.00 (0.00) 1.00 (0.00)
+Active_to_passive The artist introduced the scientist. $\rightarrow$ The scientist was introduced by the artist. Write the input sentence in passive form. 1.00 (0.00) 1.00 (0.00)
+Syntax Negation Time is finite $\rightarrow$ Time is not finite. Negate the input sentence. 0.81 (0.00) 0.82 (0.00)
+Lexical Semantics Antonyms won $\rightarrow$ lost Write a word that means the opposite of the input word. 0.70 (0.00) 0.83 (0.00)
+Synonyms alleged $\rightarrow$ supposed Write a word with a similar meaning to the input word. 0.14 (0.01) 0.13 (0.00)
+Phonetics Rhymes sing $\rightarrow$ ring Write a word that rhymes with the input word. 0.61 (0.01) 1.00 (0.00)
+Knowledge Larger_animal koala, snail $\rightarrow$ koala Write the larger of the two given animals. 0.94 (0.00) 0.94 (0.00)
 Semantics Cause_and_effect Sentence 1: The soda went flat.
@@
 Find which of the two given cause and effect sentences is the cause.
-0.97 0.92
-Common_concept guitars, pendulums, neutrinos $\rightarrow$ involve oscillations. Find a common characteristic for the given objects. 0.11 0.11
-Style Informal_to_formal Please call once you get there $\rightarrow$ Please call upon your arrival. Rephrase the sentence in formal language. 0.63 0.5
-Numerical Sum 22 10 $\rightarrow$ 32 Sum the two given numbers. 1.00 1.00
-Diff 32 22 $\rightarrow$ 10 Subtract the second number from the first. 1.00 1.00
-Num_to_Verbal 26 $\rightarrow$ twenty-six Write the number in English words. 1.00 1.00
-Multi- lingual Translation_en-de game $\rightarrow$ Spiel Translate the word into German. 0.81 0.84
-Translation_en-es game $\rightarrow$ juego Translate the word into Spanish. 0.89 0.88
-Translation_en-fr game $\rightarrow$ jeu Translate the word into French. 0.86 0.87
-GLUE Sentiment The film is small in scope, yet perfectly formed. $\rightarrow$ positive Determine whether a movie review is positive or negative. 0.89 0.90
+0.97 (0.02) 0.90 (0.02)
+Common_concept guitars, pendulums, neutrinos $\rightarrow$ involve oscillations. Find a common characteristic for the given objects. 0.11 (0.01) 0.11 (0.00)
+Style Informal_to_formal Please call once you get there $\rightarrow$ Please call upon your arrival. Rephrase the sentence in formal language. 0.63 (0.00) 0.50 (0.00)
+Numerical Sum 22 10 $\rightarrow$ 32 Sum the two given numbers. 1.00 (0.00) 1.00 (0.00)
+Diff 32 22 $\rightarrow$ 10 Subtract the second number from the first. 1.00 (0.00) 1.00 (0.00)
+Num_to_Verbal 26 $\rightarrow$ twenty-six Write the number in English words. 1.00 (0.00) 1.00 (0.00)
+Multi- lingual Translation_en-de game $\rightarrow$ Spiel Translate the word into German. 0.81 (0.00) 0.84 (0.00)
+Translation_en-es game $\rightarrow$ juego Translate the word into Spanish. 0.89 (0.00) 0.88 (0.00)
+Translation_en-fr game $\rightarrow$ jeu Translate the word into French. 0.86 (0.00) 0.87 (0.00)
+GLUE Sentiment The film is small in scope, yet perfectly formed. $\rightarrow$ positive Determine whether a movie review is positive or negative. 0.89 (0.01) 0.91 (0.00)
 Sentence_similarity Sentence 1: A man is smoking.
@@
 $\rightarrow$ 0 - definitely not Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly.
-0.00 0.21
-median score 0.89 0.90
-average score 0.78 0.79
+0.00 (0.00) 0.21 (0.00)
+median score 0.89 0.91
+average score 0.78 0.80
 # best-performing tasks 14 16
@@
 (2023).
-We tested these instructions, report their test scores, and compare them to our best test scores using ACING with Vicuna-13B as the white-box LLM.
-Appendix H Our Best Learned Instructions
+We tested these instructions, report their test scores (mean over 3 runs, standard deviation in parentheses), and compare them to our best test scores using ACING with Vicuna-13B as the white-box model.
+Appendix H Assessing Instruction Clarity: Human and Readability Analyses
+To assess the interpretability and accessibility of the instructions generated by ACING, we conduct both a human evaluation and an automated readability analysis.
+H.1 Human Evaluation
+We conducted a human annotation study across all instruction-induction tasks reported in Table˜2.
+A total of 26 participants (none of whom are paper co-authors) volunteered to independently rate the clarity, coherence, and task faithfulness of instructions produced by ACING using a 5-point Likert scale.
+The following question was used to guide ratings:
+Does the instruction clearly describe what is happening in the demonstration? Could a language model complete the task correctly by following this instruction alone?
+Each instruction was presented alongside:
+  • The task name and category
+  • An input–output demonstration pair
+  • The corresponding ACING-generated instruction
+Results.
+The results of the human evaluation are summarized below:
+  • Maximum score: 4.92
+  • Third quartile (Q3): 4.50
+  • Median (Q2): 3.90
+  • First quartile (Q1): 2.83
+  • Mean score: 3.66  (SD = 1.02)
+These results indicate that the majority of instructions are perceived as clear and well-aligned with the task by human annotators.
+H.2 Automated Readability Analysis
+To complement the human study, we applied three established readability formulas to quantify the linguistic accessibility of the 30 generated instructions in Table˜21:
+  • Flesch Reading Ease (FRE) Flesch (1948): Scores range from 0 to 100, with higher scores indicating greater ease of reading.
+  • Flesch-Kincaid Grade Level (FKG) Kincaid et al. (1975): Maps readability to U.S. school grade levels, with lower scores indicating simpler text.
+  • Coleman-Liau Index (CLI) Coleman and Liau (1975): A character-based grade-level readability metric.
+Findings.
+Readability scores across instructions show the following trends:
+  • Most instructions fall within the FRE range of 60–95, with a median 70.8, indicating that they are accessible to a general audience.
+  • FKG scores mostly range between 3 and 9, with a median 7.0, consistent with middle school to early high school reading levels.
+  • CLI scores mostly range between 2–10, with a median of 7.3.
+This aligns with the FKG analysis, indicating that instructions are generally suitable for readers with middle school to early high school reading levels.
+Overall, these results confirm that ACING generates instructions that are not only effective—according to human judgment—but also accessible, as measured by standardized readability metrics.
+The full set of readability scores is included in Table˜20.
+Instruction FRE $\uparrow$ FKG $\downarrow$ CLI $\downarrow$
+Change the input to match the output, but the output is already in the passive voice 74.3 6.9 7.00
+Take a word and change it to its opposite 94.3 2.3 2.18
+Match the input to the output, and the answer is:\$Input\$: Nature Nanotechnology, Annual Review of Biochemistry, and The Lancet Neurology \$Output\$: top journals \$Input\$: Jeans, Tops, and Suits \$Output\$: Apparel 15.6 18.5 14.53
+Input the code into a Python interpreter and observe the output 49.5 9.1 9.45
+Find the sentence that is the cause of the effect in the pair of sentences 84.5 5.2 5.43
+Make a connection between the input and output, but the connection is not clear 65.7 7.6 9.01
+Find the difference between the two numbers 66.8 5.7 10.63
+Create a function that takes a string as input and returns the first letter of the first word in the string 80.8 7.2 6.82
+Convert the input into output using the same word order and with the same meaning 67.5 7.6 8.13
+Create a program that takes two animals as input and outputs the animal that is bigger 58.4 9.1 8.09
+Input the word “year” and the output was “y e a r” 96.0 2.9 -1.35
+Flip the truth value of the statements in the input 86.7 3.7 5.60
+Convert numbers to words 75.9 3.7 7.25
+Provide a number that represents how many items are in the input 60.7 7.8 7.35
+Find the word that does not belong in each group based on the given words 95.7 3.6 5.04
+Find a word in the text that starts with the letter provided and to output that word 85.1 5.6 5.66
+Find the name of the element based on its atomic number 72.6 5.9 5.24
+Input the word that the program thought I was inputting and then output the word that program thought I was inputting 76.7 7.8 9.58
+Input a word and output the letter that corresponds to the second letter in that word 69.0 7.6 7.72
+Find a sentence pair that is probably not similar, and the output is 3 - probably 61.9 8.4 6.97
+Classify each input as positive or negative based on the assessment of the corresponding movie 33.7 12.3 13.16
+Add the suffix -s to the end of the word to make it plural 95.9 3.4 0.31
+Find the sum of the two numbers 103.0 0.6 0.69
+Input a word that is a synonym for the word that was output 83.0 4.9 2.89
+Make the AI generate a sequence of animals based on the input provided 57.0 8.5 7.80
+Provide a translation for each word in the English text into German 67.8 6.8 8.80
+Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 66.4 8.5 10.59
+Create a program that would take an English word as input and output its French equivalent 63.7 8.4 9.54
+Output the words in alphabetical order, but the output is not in alphabetical order 35.5 11.8 10.67
+Convert the input to a word that is a common English word 81.9 4.8 3.97
+Median 70.8 7.0 7.3
+Table 20: Readability analysis of the best instructions generated by ACING, measured using Flesch Reading Ease (FRE) Flesch (1948), Flesch-Kincaid Grade Level (FKG) Kincaid et al.
+(1975), and Coleman-Liau Index (CLI) Coleman and Liau (1975).
+Higher FRE and lower FKG/CLI indicate easier-to-understand instructions.
+Appendix I Our Best Learned Instructions
+In this section, we present the best-learned instructions discovered by ACING for each of the 30 instruction induction tasks.
+These instructions were generated using Vicuna-13B as the white-box model and optimized to maximize performance on the corresponding black-box evaluation.
+Table 21 showcases the resulting instructions alongside their test scores, which reflect the accuracy or task-specific metric obtained on held-out examples.
+The diversity and clarity of the instructions demonstrate ACING’s ability to synthesize task-relevant, semantically grounded prompts that elicit strong responses from black-box LLMs.
+Notably, several tasks achieve perfect scores, while others expose task-specific challenges (e.g., common_concept and synonyms), highlighting the varying difficulty across the instruction spectrum.
 Task Best instruction Test Score
@@
 rhymes Input the word that the program thought I was inputting and then output the word that program thought I was inputting 1.00
-second_word_letter Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: 0.59
+second_word_letter Input a word and output the letter that corresponds to the second letter in that word 0.91
 sentence_similarity Find a sentence pair that is probably not similar, and the output is 3 - probably 0.21
@@
 word_sorting Output the words in alphabetical order, but the output is not in alphabetical order 0.73
-word_unscrambling Convert the input to a word that is a common English word 0.62
-Table 10: The best instruction discovered by ACING for all the 30 instruction-induction tasks using with Vicuna-13B as the white-box LLM.
+word_unscrambling Convert the input to a word that is a common English word 0.63
+Table 21: The best instruction discovered by ACING for all the 30 instruction-induction tasks using with Vicuna-13B as the white-box model.
+Appendix J Use of AI Assistance.
+We used AI assistants (e.g., ChatGPT) in a limited and supporting role during the preparation of this paper.
+Specifically, we used AI tools to assist with editing text for clarity and code debugging.
+All core ideas, algorithms, experiments, results, analyses, and technical writing were fully developed and executed by the authors.
+No AI system contributed to scientific decisions, modeling choices, or interpretation of findings.